Course 311: Michaelmas Term 2001 Part II: Topics in Group Theory

D. R. Wilkins

Copyright \begin{matrix}c \\ ◯ \end{matrix} David R. Wilkins 1997–2001

Contents

2 Topics in Group Theory 2 2.1 Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 Examples of Groups . . . . . . . . . . . . . . . . . . . . . . . 3 2.3 Elementary Properties of Groups . . . . . . . . . . . . . . . . 4 2.4 Subgroups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.5 Cyclic Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.6 Cosets and Lagrange’s Theorem . . . . . . . . . . . . . . . . . 8 2.7 Normal Subgroups and Quotient Groups . . . . . . . . . . . . 10 2.8 Homomorphisms . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.9 The Isomorphism Theorems . . . . . . . . . . . . . . . . . . . 15 2.10 Direct products of groups . . . . . . . . . . . . . . . . . . . . 16 2.11 Cayley’s Theorem . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.12 Group Actions, Orbits and Stabilizers . . . . . . . . . . . . . . 17 2.13 Conjugacy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.14 Permutations and the Symmetric Groups . . . . . . . . . . . . 19 2.15 The Alternating Groups . . . . . . . . . . . . . . . . . . . . . 23 2.16 Normal Subgroups of the Symmetric Groups . . . . . . . . . . 26 2.17 Finitely Generated Abelian Groups . . . . . . . . . . . . . . . 27 2.18 The Class Equation of a Finite Group . . . . . . . . . . . . . . 30 2.19 Cauchy’s Theorem . . . . . . . . . . . . . . . . . . . . . . . . 30 2.20 The Structure of p-Groups . . . . . . . . . . . . . . . . . . . . 31 2.21 The Sylow Theorems . . . . . . . . . . . . . . . . . . . . . . . 32 2.22 Some Applications of the Sylow Theorems . . . . . . . . . . . 34 2.23 Simple Groups . . . . . . . . . . . . . . . . . . . . . . . . . . 36 2.24 Solvable Groups . . . . . . . . . . . . . . . . . . . . . . . . . . 41

2 Topics in Group Theory

2.1 Groups

A binary operation ∗ on a set G associates to elements x and y of G a third element x ∗ y of G. For example, addition and multiplication are binary operations of the set of all integers.

Definition A group G consists of a set G together with a binary operation ∗ for which the following properties are satisfied:

• (x ∗ y) ∗ z = x ∗ (y ∗ z) for all elements x, y, and z of G (the Associative Law);

• there exists an element e of G (known as the identity element of G) such that e ∗ x = x = x ∗ e, for all elements x of G;

• for each element x of G there exists an element x′ of G (known as the inverse of x) such that x ∗ x′ = e = x′ ∗ x (where e is the identity element of G).

The order |G| of a finite group G is the number of elements of G.

A group G is Abelian (or commutative) if x ∗ y = y ∗ x for all elements x and y of G.

One usually adopts multiplicative notation for groups, where the product x ∗ y of two elements x and y of a group G is denoted by xy. The inverse of an element x of G is then denoted by x^{−1}. The identity element is usually denoted by e (or by e_{G} when it is necessary to specify explicitly the group to which it belongs). Sometimes the identity element is denoted by 1. Thus, when multiplicative notation is adopted, the group axioms are written as follows:-

• (xy)z = x(yz) for all elements x, y, and z of G (the Associative Law);

• there exists an element e of G (known as the identity element of G) such that ex = x = xe, for all elements x of G;

• for each element x of G there exists an element x^{−1} of G (known as the inverse of x) such that xx^{−1} = e = x^{−1}x (where e is the identity element of G).

The group G is said to be Abelian (or commutative) if xy = yx for all elements x and y of G.

It is sometimes convenient or customary to use additive notation for certain groups. Here the group operation is denoted by +, the identity element of the group is denoted by 0, the inverse of an element x of the group is denoted by −x. By convention, additive notation is only used for Abelian groups. When expressed in additive notation the axioms for a Abelian group are as follows:

• x + y = y + x for all elements x and y of G (the Commutative Law);

• (x+y)+z = x+(y+z) for all elements x, y, and z of G (the Associative Law);

• there exists an element 0 of G (known as the identity element or zero element of G) such that 0 + x = x = x + 0, for all elements x of G;

• for each element x of G there exists an element −x of G (known as the inverse of x) such that x + (−x) = 0 = (−x) + x (where 0 is the identity element of G).

We shall usually employ multiplicative notation when discussing general properties of groups. Additive notation will be employed for certain groups (such as the set of integers with the operation of addition) where this notation is the natural one to use.

2.2 Examples of Groups

The sets of integers, rational numbers, real numbers and complex numbers are Abelian groups, where the group operation is the operation of addition.

The sets of non-zero rational numbers, non-zero real numbers and nonzero complex numbers are also Abelian groups, where the group operation is the operation of multiplication.

For each positive integer m the set ℤ_{m} of congruence classes of integers modulo m is a group, where the group operation is addition of congruence classes.

For each positive integer m the set ℤ_{m}^{∗} of congruence classes modulo m of integers coprime to m is a group, where the group operation is multiplication of congruence classes.

In particular, for each prime number p the set ℤ_{p}^{∗} of congruence classes modulo p of integers not divisible by p is a group, where the group operation is multiplication of congruence classes.

For each positive integer n the set of all nonsingular n × n matrices is a group, where the group operation is matrix multiplication. These groups are not Abelian when n ≥ 2.

Let E^{n} denote n-dimensional Euclidean space, so that E^{2} denotes the Euclidean plane, and E^{3} denotes three-dimensional Euclidean space. A geometrical figure may be represented as a subset S of E^{n}. A symmetry of S is a transformation T: E^{n} → E^{n} of E^{n} which sends straight lines to straight lines, preserves all lengths and angles, and has the property that T (S) = S. The collection of all symmetries of a geometrical figure is a group, the symmetry group of S, the group operation being that of composition of transformations.

For any natural number n greater than 2, the the dihedral group D_{2n} of order 2n is defined to be the symmetry group of a regular n-sided polygon in the Euclidean plane. It consists of rotations though an angle of 2πj/n about the centre of the polygon for j = 0, 1, 2, …, n − 1, together with the reflections in the n axes of symmetry of the polygon.

The symmetries of a rectangle that is not a square constitute a group of order 4. This group consists of the identity transformation, reflection in the axis of symmetry joining the midpoints of the two shorter sides, reflection in the axis of symmetry joining the two longer sides, and rotation though an angle of π radians (180^{∘}). If I denotes the identity transformation, A and B denote the reflections in the two axes of symmetry, and C denotes the rotation through π radians then A^{2} = B^{2} = C^{2} = I, AB = BA = C, AC = CA = B and BC = CB = A. This group is Abelian: it is often referred to as the Klein 4-group (or, in German, Kleinsche Viergruppe).

The symmetries of a regular tetrahedron in 3-dimensional space constitute a group. Any permutation of the vertices of the tetrahedron can be effected by an appropriate symmetry of the tetrahedron. Moreover each symmetry is completely determined by the permutation of the vertices which it induces. Therefore the group of symmetries of a regular tetrahedron is of order 24, since there are 24 permutations of a set with four elements. It turns out that this group is non-Abelian.

2.3 Elementary Properties of Groups

In what follows, we describe basic properties of a group G, using multiplicative notation and denoting the identity element of the group by the letter e.

Lemma 2.1 A group G has exactly one identity element e satisfying ex = x = xe for all x ∈ G.

Proof Suppose that f is an element of G with the property that fx = x for all elements x of G. Then in particular f = fe = e. Similarly one can show that e is the only element of G satisfying xe = x for all elements x of G.

Lemma 2.2 An element x of a group G has exactly one inverse x^{−1}.

Proof We know from the axioms that the group G contains at least one element x^{−1} which satisfies xx^{−1} = e and x^{−1}x = e. If z is any element of G which satisfies xz = e then z = ez = (x^{−1}x)z = x^{−1}(xz) = x^{−1}e = x^{−1}. Similarly if w is any element of G which satisfies wx = e then w = x^{−1}. In particular we conclude that the inverse x^{−1} of x is uniquely determined, as required.

Lemma 2.3 Let x and y be elements of a group G. Then (xy)^{−1} = y^{−1}x^{−1}.

Proof It follows from the group axioms that

(xy)(y^{−1}x^{−1}) = x(y(y^{−1}x^{−1})) = x((yy^{−1})x^{−1}) = x(ex^{−1}) = xx^{−1} = e.

Similarly (y^{−1}x^{−1})(xy) = e, and thus y^{−1}x^{−1} is the inverse of xy, as required.

Note in particular that (x^{−1})^{−1} = x for all elements x of a group G, since x has the properties that characterize the inverse of the inverse x^{−1} of x.

Given an element x of a group G, we define x^{n} for each positive integer n by the requirement that x^{1} = x and x^{n} = x^{n−1}x for all n > 1. We also define x^{0} = e, where e is the identity element of the group, and we define x^{−n} to be the inverse of x^{n} for all positive integers n.

Theorem 2.4 Let x be an element of a group G. Then x^{m+n} = x^{m}x^{n} and x^{mn} = (x^{m})^{n} for all integers m and n.

Proof The identity x^{m+n} = x^{m}x^{n} clearly holds when m = 0 and when n = 0. The identity x^{m+n} = x^{m}x^{n} can be proved for all positive integers m and n by induction on n. The identity when m and n are both negative then follows from the identity x^{−m−n} = x^{−n}x^{−m} on taking inverses. The result when m and n have opposite signs can easily be deduced from that where m and n both have the same sign.

The identity x^{mn} = (x^{m})^{n} follows immediately from the definitions when n = 0, 1 or −1. The result when n is positive can be proved by induction on n. The result when n is negative can then be obtained on taking inverses.

If additive notation is employed for an Abelian group then the notation ‘x^{n}’ is replaced by ‘nx’ for all integers n and elements x of the group. The analogue of Theorem 2.4 then states that (m + n)x = mx + nx and (mn)x = m(n(x)) for all integers m and n.

The associative law may be generalized to products of four or more elements of a group.

Example Given four elements x_{1}, x_{2}, x_{3} and x_{4} of a group, the products

((x_{1}x_{2})x_{3})x_{4}, (x_{1}x_{2})(x_{3}x_{4}), (x_{1}(x_{2}x_{3}))x_{4}, x_{1}((x_{2}x_{3})x_{4}), x_{1}(x_{2}(x_{3}x_{4}))

all have the same value. (Note that x_{1}x_{2}x_{3}x_{4} is by definition the value of the first of these expressions.)

Two expressions, each specifying a finite product of elements of a group G, determine the same element of G if the same elements of G occur in both expressions, and in the same order. This result can be proved by induction on the number of elements of G making up such a product.

2.4 Subgroups

Definition Let G be a group, and let H be a subset of G. We say that H is a subgroup of G if the following conditions are satisfied:

• the identity element of G is an element of H;

• the product of any two elements of H is itself an element of H;

• the inverse of any element of H is itself an element of H.

Lemma 2.5 Let x be an element of a group G. Then the set of all elements of G that are of the form x^{n} for some integer n is a subgroup of G.

Proof Let H = {x^{n} : n ∈ ℤ}. Then the identity element belongs to H, since it is equal to x^{0}. The product of two elements of H is itself an element of H, since x^{m}x^{n} = x^{m+n} for all integers m and n (see Theorem 2.4). Also the inverse of an element of H is itself an element of H since (x^{n})^{−1} = x^{−n} for all integers n. Thus H is a subgroup of G, as required.

Definition Let x be an element of a group G. The order of x is the smallest positive integer n for which x^{n} = e. The subgroup generated by x is the subgroup consisting of all elements of G that are of the form x^{n} for some integer n.

Lemma 2.6 Let H and K be subgroups of a group G. Then H ∩ K is also a subgroup of G.

Proof The identity element of G belongs to H ∩ K since it belongs to the subgroups H and K. If x and y are elements of H ∩ K then xy is an element of H (since x and y are elements of H), and xy is an element of K, and therefore xy is an element of H ∩ K. Also the inverse x^{−1} of an element x of H ∩ K belongs to H and to K and thus belongs to H ∩ K, as required.

More generally, the intersection of any collection of subgroups of a given group is itself a subgroup of that group.

2.5 Cyclic Groups

Definition A group G is said to be cyclic, with generator x, if every element of G is of the form x^{n} for some integer n.

Example The group ℤ of integers under addition is a cyclic group, generated by 1.

Example Let n be a positive integer. The set ℤ_{n} of congruence classes of integers modulo n is a cyclic group of order n with respect to the operation of addition.

Example The group of all rotations of the plane about the origin through an integer multiple of 2π/n radians is a cyclic group of order n for all integers n. This group is generated by an anticlockwise rotation through an angle of 2π/n radians.

Lemma 2.7 Let G be a finite cyclic group with generator x, and let j and k be integers. Then x^{j} = x^{k} if and only if j − k is divisible by the order of the group.

Proof First we show that x^{m} = e for some strictly positive integer m, where e is the identity element of G. Now x^{j} = x^{k} for some integers j and k with j < k, since G is finite. Let m = k − j. Then m > 0 and x^{m} = x^{k}(x^{j})^{−1} = e.

Let n be the smallest strictly positive integer for which x^{n} = e. Now any integer i can be expressed in the form i = qn + r, where q and r are integers and 0 ≤ r < n. (Thus q is the greatest integer for which qn ≤ i.) Then x^{i} = (x^{n})^{q}x^{r} = x^{r} (since x^{n} = e). Now the choice of n ensures that x^{r} ≠ e if 0 < r < n. It follows that an integer i satisfies x^{i} = e if and only if n divides i.

Let j and k be integers. Now x^{j} = x^{k} if and only if x^{j−k} = e, since x^{j−k} = x^{j}(x^{k})^{−1}. It follows that x^{j} = x^{k} if and only if j − k is divisible by n. Moreover n is the order of the group G, since each element of G is equal to one of the elements x^{i} with 0 ≤ i < n and these elements are distinct.

We now classify all subgroups of a cyclic group G. Let x be a generator of G. Given a subgroup H of G with more than one element, let m be the smallest strictly positive integer for which x^{m} ∈ H. Suppose that x^{i} ∈ H for some integer i. Now i can be expressed in the form i = qm + r, where q and r are integers and 0 ≤ r < m. (Thus q is the greatest integer for which qm ≤ i.) But then x^{r} = x^{i−qm} = x^{i}(x^{m})^{−q}, where x^{i} ∈ H and x^{m} ∈ H, and therefore x^{r} ∈ H. The choice of m now ensures that r = 0, and hence i = qm. Thus x^{i} ∈ H if and only if i is some integer multiple of m. This shows that H is the cyclic group generated by x^{m}, where m is the smallest strictly positive integer for which x^{m} ∈ H.

Let us consider the case when the cyclic group G is finite. Let s be the order of G. Then x^{s} = e, and hence x^{s} belongs to the subgroup H. It follows that s must be some integer multiple of m, where m is the smallest strictly positive integer for which x^{m} ∈ H. Thus the subgroups of a finite cyclic group G with generator g are the trivial subgroup {e} and the cyclic subgroups generated by x^{m} for each divisor m of the order of G.

Consider now the case when the cyclic group G is infinite. For each positive integer m, the element x^{m} generates a subgroup of G, and moreover m is the smallest strictly positive integer for which x^{m} belongs to that subgroup. Thus if G is an infinite cyclic group with generator x then the subgroups of G are the trivial subgroup {e} and the cyclic subgroups generated by x^{m} for each positive integer m.

We have thus classified all subgroups of a cyclic group. In particular we see that any subgroup of a cyclic group is itself a cyclic group.

2.6 Cosets and Lagrange’s Theorem

Definition Let H be a subgroup of a group G. A left coset of H in G is a subset of G that is of the form xH, where x ∈ G and

xH = {y ∈ G : y = xh for some h ∈ H}.

Similarly a right coset of H in G is a subset of G that is of the form Hx, where x ∈ G and

Hx = {y ∈ G : y = hx for some h ∈ H}.

Note that a subgroup H of a group G is itself a left coset of H in G.

Lemma 2.8 Let H be a subgroup of a group G. Then the left cosets of H in G have the following properties:—

(i) x ∈ xH for all x ∈ G;

(ii) if x and y are elements of G, and if y = xa for some a ∈ H, then xH = yH;

(iii) if x and y are elements of G, and if xH ∩ yH is non-empty then xH = yH.

Proof Let x ∈ G. Then x = xe, where e is the identity element of G. But e ∈ H. It follows that x ∈ xH. This proves (i).

Let x and y be elements of G, where y = xa for some a ∈ H. Then yh = x(ah) and xh = y(a^{−1}h) for all h ∈ H. Moreover ah ∈ H and a^{−1}h ∈ H for all h ∈ H, since H is a subgroup of G. It follows that yH ⊂ xH and xH ⊂ yH, and hence xH = yH. This proves (ii).

Finally suppose that xH ∩ yH is non-empty for some elements x and y of G. Let z be an element of xH ∩ yH. Then z = xa for some a ∈ H, and z = yb for some b ∈ H. It follows from (ii) that zH = xH and zH = yH. Therefore xH = yH. This proves (iii).

Lemma 2.9 Let H be a finite subgroup of a group G. Then each left coset of H in G has the same number of elements as H.

Proof Let H = {h_{1}, h_{2}, …, h_{m}}, where h_{1}, h_{2}, …, h_{m} are distinct, and let x be an element of G. Then the left coset xH consists of the elements xh_{j} for j = 1, 2, …, m. Suppose that j and k are integers between 1 and m for which xh_{j} = xh_{k}. Then h_{j} = x^{−1}(xh_{j}) = x^{−1}(xh_{k}) = h_{k}, and thus j = k, since h_{1}, h_{2}, …, h_{m} are distinct. It follows that the elements xh_{1}, xh_{2}, …, xh_{m} are distinct. We conclude that the subgroup H and the left coset xH both have m elements, as required.

Theorem 2.10 (Lagrange’s Theorem) Let G be a finite group, and let H be a subgroup of G. Then the order of H divides the order of G.

Proof Each element of G belongs to at least one left coset of H in G, and no element can belong to two distinct left cosets of H in G (see Lemma 2.8). Therefore every element of G belongs to exactly one left coset of H. Moreover each left coset of H contains |H| elements (Lemma 2.9). Therefore |G| = n|H|, where n is the number of left cosets of H in G. The result follows.

Definition Let H be a subgroup of a group G. If the number of left cosets of H in G is finite then the number of such cosets is referred to as the index of H in G, denoted by [G: H].

The proof of Lagrange’s Theorem shows that the index [G: H] of a subgroup H of a finite group G is given by [G: H] = |G|/|H|.

Corollary 2.11 Let x be an element of a finite group G. Then the order of x divides the order of G.

Proof Let H be the set of all elements of G that are of the form x^{n} for some integer n. Then H is a subgroup of G (see Lemma 2.5), and the order of H is the order of x. But the order of H divides G by Lagrange’s Theorem (Theorem 2.10). The result follows.

Corollary 2.12 Any finite group of prime order is cyclic.

Proof Let G be a group of prime order, and let x be some element of G that is not the identity element. Then the order of x is greater than one and divides the order of G. But then the order of x must be equal to the order of G, since the latter is a prime number. Thus G is a cyclic group generated by x, as required.

2.7 Normal Subgroups and Quotient Groups

Let A and B be subsets of a group G. The product AB of the sets A and B is defined by

AB = {xy : x ∈ A and y ∈ B}.

We denote {x}A and A{x} by xA and Ax, for all elements x of G and subsets A of G. The Associative Law for multiplication of elements of G ensures that (AB)C = A(BC) for all subsets A, B and C of G. We can therefore use the notation ABC to denote the products (AB)C and A(BC); and we can use analogous notation to denote the product of four or more subsets of G.

If A, B and C are subsets of a group G, and if A ⊂ B then clearly AC ⊂ BC and CA ⊂ CB.

Note that if H is a subgroup of the group G and if x is an element of G then xH is the left coset of H in G that contains the element x. Similarly Hx is the right coset of H in G that contains the element x.

If H is a subgroup of G then HH = H. Indeed HH ⊂ H, since the product of two elements of a subgroup H is itself an element of H. Also H ⊂ HH since h = eh for any element h of H, where e, the identity element of G, belongs to H.

Definition A subgroup N of a group G is said to be a normal subgroup of G if xnx^{−1} ∈ N for all n ∈ N and x ∈ G.

The notation ‘N ◁ G’ signifies ‘N is a normal subgroup of G’.

Definition A non-trivial group G is said to be simple if the only normal subgroups of G are the whole of G and the trivial subgroup {e} whose only element is the identity element e of G.

Lemma 2.13 Every subgroup of an Abelian group is a normal subgroup.

Proof Let N be a subgroup of an Abelian group G. Then

xnx^{−1} = (xn)x^{−1} = (nx)x^{−1} = n(xx^{−1}) = ne = n

for all n ∈ N and x ∈ G, where e is the identity element of G. The result follows.

Example Let S_{3} be the group of permutations of the set {1, 2, 3}, and let H be the subgroup of S_{3} consisting of the identity permutation and the transposition (1 2). Then H is not normal in G, since (2 3)^{−1}(1 2)(2 3) = (2 3)(1 2)(2 3) = (1 3) and (1 3) does not belong to the subgroup H.

Proposition 2.14 A subgroup N of a group G is a normal subgroup of G if and only if xNx^{−1} = N for all elements x of G.

Proof Suppose that N is a normal subgroup of G. Let x be an element of G. Then xNx^{−1} ⊂ N. (This follows directly from the definition of a normal subgroup.) On replacing x by x^{−1} we see also that x^{−1}Nx ⊂ N, and thus N = x(x^{−1}Nx)x^{−1} ⊂ xNx^{−1}. Thus each of the sets N and xNx^{−1} is contained in the other, and therefore xNx^{−1} = N.

Conversely if N is a subgroup of G with the property that xNx^{−1} = N for all x ∈ G, then it follows immediately from the definition of a normal subgroup that N is a normal subgroup of G.

Corollary 2.15 A subgroup N of a group G is a normal subgroup of G if and only if xN = Nx for all elements x of G.

Proof Let N be a subgroup of G, and let x be an element of G. If xNx^{−1} = N then xN = (xNx^{−1})x = Nx. Conversely if xN = Nx then xNx^{−1} = Nxx^{−1} = Ne = N, where e is the identity element of G. Thus xN = Nx if and only if xNx^{−1} = N. It follows from Proposition 2.14 that a subgroup N of G is normal if and only if xN = Nx for all elements x of G, as required.

Let N be a normal subgroup of G. Corollary 2.15 shows that a subset of G is a left coset of N in G if and only if it is a right coset of N in G. We may therefore refer to the left and right cosets of a normal subgroup N as cosets of N in G (since it is not in this case necessary to distinguish between left and right cosets).

Lemma 2.16 Let N be a normal subgroup of a group G and let x and y be elements of G. Then (xN)(yN) = (xy)N.

Proof If N is a normal subgroup of G then Ny = yN, and therefore (xN)(yN) = x(Ny)N = x(yN)N = (xy)(NN). But NN = N, since N is a subgroup of G. Therefore (xN)(yN) = (xy)N, as required.

Proposition 2.17 Let G be a group, and let N be a normal subgroup of G. Then the set of all cosets of N in G is a group under the operation of multiplication. The identity element of this group is N itself, and the inverse of a coset xN is the coset x^{−1}N for any element x of G.

Proof Let x, y and z be any elements of G. Then the product of the cosets xN and yN is the coset (xy)N. The subgroup N is itself a coset of N in G, since N = eN. Moreover

(xN)N = (xN)(eN) = (xe)N = xN,
N(xN) = (eN)(xN) = (ex)N = xN,
(xN)(x^{−1}N) = (xx^{−1})N = N,
(x^{−1}N)(xN) = (x^{−1}x)N = N.

for all elements x of G. Thus the group axioms are satisfied.

Definition Let N be a normal subgroup of a group G. The quotient group G/N is defined to be the group of cosets of N in G under the operation of multiplication.

Example Consider the dihedral group D_{8} of order 8, which we represent as the group of symmetries of a square in the plane with corners at the points whose Cartesian co-ordinates are (1, 1), (−1, 1), (−1, −1) and (1, −1). Then

D_{8} = {I, R, R^{2}, R^{3}, T_{1}, T_{2}, T_{3}, T_{4}},

where I denotes the identity transformation, R denotes an anticlockwise rotation about the origin through a right angle, and T_{1}, T_{2}, T_{3} and T_{4} denote the reflections in the lines y = 0, x = y, x = 0 and x = −y respectively. Let N = {I, R^{2}}. Then N is a subgroup of D_{8}. The left cosets of N in D_{8} are N, A, B and C, where

A = {R, R^{3}},  B = {T_{1}, T_{3}},  C = {T_{2}, T_{4}}.

Moreover N, A, B and C are also the right cosets of N in D_{8}, and thus N is a normal subgroup of D_{8}. On multiplying the cosets A, B and C with one another we find that AB = BA = C, AC = CA = B and BC = CB = A. The quotient group D_{8}/N consists of the set {N, A, B, C}, with the group operation just described.

2.8 Homomorphisms

Definition A homomorphism θ: G → K from a group G to a group K is a function with the property that θ(g_{1} ∗ g_{2}) = θ(g_{1}) ∗ θ(g_{2}) for all g_{1}, g_{2} ∈ G, where ∗ denotes the group operation on G and on K.

Example Let q be an integer. The function from the group ℤ of integers to itself that sends each integer n to qn is a homomorphism.

Example Let x be an element of a group G. The function that sends each integer n to the element x^{n} is a homomorphism from the group ℤ of integers to G, since x^{m+n} = x^{m}x^{n} for all integers m and n (Theorem 2.4).

Lemma 2.18 Let θ: G → K be a homomorphism. Then θ(e_{G}) = e_{K}, where e_{G} and e_{K} denote the identity elements of the groups G and K. Also θ(x^{−1}) = θ(x)^{−1} for all elements x of G.

Proof Let z = θ(e_{G}). Then z^{2} = θ(e_{G})θ(e_{G}) = θ(e_{G}e_{G}) = θ(e_{G}) = z. The result that θ(e_{G}) = e_{K} now follows from the fact that an element z of K satisfies z^{2} = z if and only if z is the identity element of K.

Let x be an element of G. The element θ(x^{−1}) satisfies θ(x)θ(x^{−1}) = θ(xx^{−1}) = θ(e_{G}) = e_{K}, and similarly θ(x^{−1})θ(x) = e_{K}. The uniqueness of the inverse of θ(x) now ensures that θ(x^{−1}) = θ(x)^{−1}.

An isomorphism θ: G → K between groups G and K is a homomorphism that is also a bijection mapping G onto K. Two groups G and K are isomorphic if there exists an isomorphism mapping G onto K.

Example Let D_{6} be the group of symmetries of an equilateral triangle in the plane with vertices A, B and C, and let S_{3} be the group of permutations of the set {A, B, C}. The function which sends a symmetry of the triangle to the corresponding permutation of its vertices is an isomorphism between the dihedral group D_{6} of order 6 and the symmetric group S_{3}.

Example Let ℝ be the group of real numbers with the operation of addition, and let ℝ^{+} be the group of strictly positive real numbers with the operation of multiplication. The function exp: ℝ → ℝ^{+} that sends each real number x to the positive real number e^{x} is an isomorphism: it is both a homomorphism of groups and a bijection. The inverse of this isomorphism is the function log: ℝ^{+} → ℝ that sends each strictly positive real number to its natural logarithm.

Here is some further terminology regarding homomorphisms:

• A monomorphism is an injective homomorphism.

• An epimorphism is a surjective homomorphism.

• An endomorphism is a homomorphism mapping a group into itself.

• An automorphism is an isomorphism mapping a group onto itself.

Definition The kernel ker θ of the homomorphism θ: G → K is the set of all elements of G that are mapped by θ onto the identity element of K.

Example Let the group operation on the set {+1, −1} be multiplication, and let θ: ℤ → {+1, −1} be the homomorphism that sends each integer n to (−1)^{n}. Then the kernel of the homomorphism θ is the subgroup of ℤ consisting of all even numbers.

Lemma 2.19 Let G and K be groups, and let θ: G → K be a homomorphism from G to K. Then the kernel ker θ of θ is a normal subgroup of G.

Proof Let x and y be elements of ker θ. Then θ(x) = e_{K} and θ(y) = e_{K}, where e_{K} denotes the identity element of K. But then θ(xy) = θ(x)θ(y) = e_{K}e_{K} = e_{K}, and thus xy belongs to ker θ. Also θ(x^{−1}) = θ(x)^{−1} = e_{K}^{−1} = e_{K}, and thus x^{−1} belongs to ker θ. We conclude that ker θ is a subgroup of K. Moreover ker θ is a normal subgroup of G, for if g ∈ G and x ∈ ker θ then

θ(gxg^{−1}) = θ(g)θ(x)θ(g)^{−1} = θ(g)θ(g^{−1}) = e_{K}.

If N is a normal subgroup of some group G then N is the kernel of the quotient homomorphism θ: G → G/N that sends g ∈ G to the coset gN. It follows therefore that a subset of a group G is a normal subgroup of G if and only if it is the kernel of some homomorphism.

Proposition 2.20 Let G and K be groups, let θ: G → K be a homomorphism from G to K, and let N be a normal subgroup of G. Suppose that N ⊂ ker θ. Then the homomorphism θ: G → K induces a homomorphism θ̂: G/N → K sending gN ∈ G/N to θ(g). Moreover θ̂: G/N → K is injective if and only if N = ker θ.

Proof Let x and y be elements of G. Now xN = yN if and only if x^{−1}y ∈ N. Also θ(x) = θ(y) if and only if x^{−1}y ∈ ker θ. Thus if N ⊂ ker θ then θ(x) = θ(y) whenever xN = yN, and thus θ: G → K induces a well-defined function θ̂: G/N → K sending xN ∈ G/N to θ(x). This function is a homomorphism, since θ̂((xN)(yN)) = θ̂(xyN) = θ(xy) = θ(x)θ(y) = θ̂(xN)θ̂(yN).

Suppose now that N = ker θ. Then θ(x) = θ(y) if and only if xN = yN. Thus the homomorphism θ̂: G/N → K is injective. Conversely if θ̂: G/N → K is injective then N must be the kernel of θ, as required.

Corollary 2.21 Let G and K be groups, and let θ: G → K be a homomorphism. Then θ(G) ≅ G/ ker θ.

2.9 The Isomorphism Theorems

Lemma 2.22 Let G be a group, let H be a subgroup of G, and let N be a normal subgroup of G. Then the set HN is a subgroup of G, where HN = {hn : h ∈ H and n ∈ N}.

Proof The set HN clearly contains the identity element of G. Let x and y be elements of HN. We must show that xy and x^{−1} belong to HN. Now x = hu and y = kv for some elements h and k of H and for some elements u and v of N. Then xy = (hk)(k^{−1}ukv). But k^{−1}uk ∈ N, since N is normal. It follows that k^{−1}ukv ∈ N, since N is a subgroup and k^{−1}ukv is the product of the elements k^{−1}uk and v of N. Also hk ∈ H. It follows that xy ∈ HN.

We must also show that x^{−1} ∈ HN. Now x^{−1} = u^{−1}h^{−1} = h^{−1}(hu^{−1}h^{−1}). Also h^{−1} ∈ H, since H is a subgroup of G, and hu^{−1}h^{−1} ∈ N, since N is a normal subgroup of G. It follows that x^{−1} ∈ HN, and thus HN is a subgroup of G, as required.

Theorem 2.23 (First Isomorphism Theorem) Let G be a group, let H be a subgroup of G, and let N be a normal subgroup of G. Then

\frac{HN}{N} ≅ \frac{H}{N ∩ H} .

Proof Every element of HN/N is a coset of N that is of the form hN for some h ∈ H. Thus if φ(h) = hN for all h ∈ H then φ: H → HN/N is a surjective homomorphism, and ker φ = N ∩ H. But φ(H) ≅ H/ ker φ (Corollary 2.21). Therefore HN/N ≅ H/(N ∩ H) as required.

Theorem 2.24 (Second Isomorphism Theorem) Let M and N be normal subgroups of a group G, where M ⊂ N. Then

\frac{G}{N} ≅ \frac{G/M}{N/M} .

Proof There is a well-defined homomorphism θ: G/M → G/N that sends gM to gN for all g ∈ G. Moreover the homomorphism θ is surjective, and ker θ = N/M. But θ(G/M) ≅ (G/M)/ ker θ (Corollary 2.21). Therefore G/N is isomorphic to (G/M) / (N/M), as required.

2.10 Direct products of groups

Let G_{1}, G_{2}, …, G_{n} be groups, and let G be the Cartesian product G_{1} × G_{2} × ⋯ × G_{n} of G_{1}, G_{2}, …, G_{n} (when the latter are regarded as sets). Then the elements of G are n-tuples (x_{1}, x_{2}, …, x_{n}) where x_{i} ∈ G_{i} for i = 1, 2, …, n. We can multiply two elements of G as follows:

(x_{1}, x_{2}, …, x_{n})(y_{1}, y_{2}, …, y_{n}) = (x_{1}y_{1}, x_{2}y_{2}, …, x_{n}y_{n}).

One can readily verify that G is a group with respect to this binary operation: multiplication is associative; the identity element of the group is (e_{1}, e_{2}, …, e_{n}), where e_{i} is the identity element of G_{i} for each i; and the inverse of an element (x_{1}, x_{2}, …, x_{n}) of G is (x_{1}^{−1}, x_{2}^{−1}, …, x_{n}^{−1}). We say that the group G is the direct product of the groups G_{1}, G_{2}, …, G_{n}: this direct product is (not surprisingly) denoted by G_{1} × G_{2} × ⋯ × G_{n}.

Example Let C_{2} and C_{3} be cyclic groups of orders 2 and 3 respectively. Then C_{2} × C_{3} is a cyclic group of order 6, and C_{2} × C_{2} is isomorphic to the Klein 4-group that consists of the symmetries of a rectangle in the Euclidean plane that is not a square.

Let us first consider C_{2} × C_{3}. Let x and y be generators of C_{2} and C_{3} respectively, and let e and e′ denote the identity elements of C_{2} and C_{3}. Thus C_{2} = {e, x} and C_{3} = {e′, y, y^{2}}, where x^{2} = e and y^{3} = e′. The elements of C_{2} × C_{3} are

(e, e′), (e, y), (e, y^{2}), (x, e′), (x, y), (x, y^{2}).

Let z = (x, y). On computing the powers of z we find that

z^{2} = (e, y^{2}), z^{3} = (x, e′), z^{4} = (e, y), z^{5} = (x, y^{2}), z^{6} = (e, e′).

Thus 6 is the smallest positive integer n for which z^{n} is equal to the identity element (e, e′) of the group. We deduce that the group C_{2} × C_{3} (which is a group of order 6) must be a cyclic group generated by the element z.

Next consider C_{2} × C_{2}. This has four elements I, A, B and C, where I = (e, e), A = (e, x), B = (x, e) and C = (x, x). If we calculate the Cayley table for the group, we discover that it is that of the Klein 4-group.

2.11 Cayley’s Theorem

Theorem 2.25 (Cayley’s Theorem) Let G be a group of order n. Then G is isomorphic to a subgroup of the group S_{n} of permutations of a set of n elements.

Proof For each element x of G, let σ_{x}: G → G be the function defined such that σ_{x}(g) = xg for all g ∈ G. Now

σ_{x^{−1}}(σ_{x}(g)) = x^{−1}(xg) = (x^{−1}x)g = g

and

σ_{x}(σ_{x^{−1}}(g)) = x(x^{−1}g) = (xx^{−1})g = g

for all g ∈ G. It follows that, for any x ∈ G, the function σ_{x}: G → G is a bijection whose inverse is σ_{x^{−1}}. It follows that σ_{x} is a permutation of G for all x ∈ G, and thus the function sending an element x of G to the permutation σ_{x} is a function from G to the group of permutations of G. This function is a homomorphism. Indeed σ_{xy} = σ_{x} ∘ σ_{y} since σ_{xy}(g) = (xy)g = x(yg) = σ_{x}(σ_{y}(g)) for all g ∈ G. The homomorphism sending x ∈ G to σ_{x} is be injective, for if σ_{x} is the identity permutation then xg = g for all g ∈ G, and hence x is the identity element of G. It follows that G is isomorphic to the image of the homomorphism. This image is a subgroup {σ_{x} : x ∈ G} of the group of permutations of G. The result follows.

2.12 Group Actions, Orbits and Stabilizers

Definition A left action of a group G on a set X associates to each g ∈ G and x ∈ X an element g.x of X in such a way that g.(h.x) = (gh).x and 1.x = x for all g, h ∈ G and x ∈ X, where 1 denotes the identity element of G.

Given a left action of a group G on a set X, the orbit of an element x of X is the subset {g.x : g ∈ G} of X, and the stabilizer of x is the subgroup {g ∈ G : g.x = x} of G.

Lemma 2.26 Let G be a finite group which acts on a set X on the left. Then the orbit of an element x of X contains [G: H] elements, where [G: H] is the index of the stabilizer H of x in G.

Proof There is a well-defined function θ: G/H → X defined on the set G/H of left cosets of H in G which sends gH to g.x for all g ∈ G. Moreover this function is injective, and its image is the orbit of x. The result follows.

2.13 Conjugacy

Definition Two elements h and k of a group G are said to be conjugate if k = ghg^{−1} for some g ∈ G.

One can readily verify that the relation of conjugacy is reflexive, symmetric and transitive and is thus an equivalence relation on a group G. The equivalence classes determined by this relation are referred to as the conjugacy classes of G. A group G is the disjoint union of its conjugacy classes. Moreover the conjugacy class of the identity element of G contains no other element of G.

A group G is Abelian if and only if all its conjugacy classes contain exactly one element of the group G.

Definition Let G be a group. The centralizer C(h) of an element h of G is the subgroup of G defined by C(h) = {g ∈ G : gh = hg}.

Lemma 2.27 Let G be a finite group, and let h ∈ G. Then the number of elements in the conjugacy class of h is equal to the index [G: C(h)] of the centralizer C(h) of h in G.

Proof There is a well-defined function f: G/C(h) → G, defined on the set G/C(h) of left cosets of C(h) in G, which sends the coset gC(h) to ghg^{−1} for all g ∈ G. This function is injective, and its image is the conjugacy class of h. The result follows.

Let H be a subgroup of a group G. One can easily verify that gHg^{−1} is also a subgroup of G for all g ∈ G, where gHg^{−1} = {ghg^{−1} : h ∈ H}.

Definition Two subgroups H and K of a group G are said to be conjugate if K = gHg^{−1} for some g ∈ G.

The relation of conjugacy is an equivalence relation on the collection of subgroups of a given group G.

2.14 Permutations and the Symmetric Groups

A permutation of a set S is a bijective function p: S → S from S to itself.

The identity permutation of a set S is the permutation that fixes every element of S.

Permutations of a finite set S are conveniently represented in a two row form

( \begin{matrix}x_{1} & x_{2} & … & x_{n} \\ p(x_{1}) & p(x_{2}) & … & p(x_{n}) \end{matrix} ) ,

where x_{1}, x_{2}, …, x_{n} are the elements of the set S and p(x_{1}), p(x_{2}), …, p(x_{n}) are the images of these elements under the permutation p being represented. Thus for example

( \begin{matrix}1 & 2 & 3 \\ 2 & 3 & 1 \end{matrix} )

represents the permutation of the set {1, 2, 3} that sends 1 to 2, sends 2 to 3, and sends 3 to 1.

Example There are two permutations of a set {a, b} with two elements.

These are the identity permutation ( \begin{matrix}a & b \\ a & b \end{matrix} ) and the transposition ( \begin{matrix}a & b \\ b & a \end{matrix} )

that interchanges the elements a and b.

Example There are six permutations of a set {a, b, c} with three elements. These are

( \begin{matrix}a & b & c \\ a & b & c \end{matrix} ) , ( \begin{matrix}a & b & c \\ a & c & b \end{matrix} ) , ( \begin{matrix}a & b & c \\ b & a & c \end{matrix} ) ,

( \begin{matrix}a & b & c \\ b & c & a \end{matrix} ) , ( \begin{matrix}a & b & c \\ c & a & b \end{matrix} ) , ( \begin{matrix}a & b & c \\ c & b & a \end{matrix} ) .

Let S be a set. Then the composition of any two permutations of S is itself a permutation of S (since the composition of two bijections is a bijection). Also any permutation p of S has a well-defined inverse p^{−1}. (This follows from the fact that the inverse of a bijection is itself a bijection.) Composition of permutations is associative: (p ∘ q) ∘ r = p ∘ (q ∘ r) for all permutations p, q and r of S. (This can be verified by noting that ((p∘q)∘r)(x) = p(q(r(x))) = (p ∘ (q ∘ r))(x) for all elements x of S.) It follows from this that the set of all permutations of a set S is a group, where the group operation is composition of permutations.

Definition For each natural number n, the symmetric group Σ_{n} is the group of permutations of the set {1, 2, …, n}.

Let S be a set, and let a_{1}, a_{2}, …, a_{n} be distinct elements of S. We denote by (a_{1} a_{2} ⋯ a_{n}) the permutation of S that sends a_{i} to a_{i+1} for i = 1, 2, …, n − 1, sends a_{n} to a_{1}, and fixes all other elements of S. Such a permutation is called a cycle of order n, or n-cycle. A cycle of length 2 is also called a transposition.

(Note that evaluating a composition of cycles, we shall compose them from right to left, in accordance with standard practice when composing functions.)

Example There are 24 permutations of a set {a, b, c, d} with exactly four elements. These are the following: the identity permutation that fixes every element of the set; the six transpositions (a b), (a c), (a d), (b c), (b d) and (c d); the eight 3-cycles (b c d), (b d c), (a c d), (a d c), (a b d), (a d b), (a b c) and (a c b); the six 4-cycles (a b c d), (a b d c), (a c b d), (a c d b), (a d b c) and (a d c b); and three further permutations (a b)(c d), (a c)(b d) and (a d)(b c).

Two cycles (a_{1} a_{2} ⋯ a_{m}) and (b_{1} b_{2} ⋯ b_{n}) are said to be disjoint when the elements a_{1}, a_{2}, …, a_{m} and b_{1}, b_{2}, …, b_{n} are distinct (i.e., no pair of these elements coincide).

It is easy to see that if (a_{1} a_{2} ⋯ a_{m}) and (b_{1} b_{2} ⋯ b_{n}) are disjoint cycles then

(a_{1} a_{2} ⋯ a_{m})(b_{1} b_{2} ⋯ b_{n}) = (b_{1} b_{2} ⋯ b_{n})(a_{1} a_{2} ⋯ a_{m}).

Proposition 2.28 Any permutation of a finite set S is the identity permutation, a cycle, or a composition of two or more disjoint cycles.

Proof We prove the result by induction on the number of elements in the set S. The result is trivially true if S has only one element, since in this case the only permutation of S is the identity permutation. Suppose that the result is known to be true for all permutations of sets with fewer than k elements. We show that the result then holds for all permutations of sets with k elements.

Let S be a set with k elements and let p be a permutation of S. Choose an element a_{1} of S, and let elements a_{2}, a_{3}, a_{4}, … of S be defined by the requirement that p(a_{i}) = a_{i+1} for all positive integers i. Let n be the largest positive integer for which the elements a_{1}, a_{2}, …, a_{n} of S are distinct. We claim that p(a_{n}) = a_{1}.

Now the choice of n ensures that the elements a_{1}, a_{2}, …, a_{n}, a_{n+1} are not distinct. Therefore a_{n+1} = a_{j} for some positive integer j between 1 and n. If j were greater than one then we would have a_{j} = p(a_{j−1}) and a_{j} = p(a_{n}), which is impossible since if p is a permutation of S then exactly one element of S must be sent to a_{j} by p. Therefore j = 1, and thus p(a_{n}) = a_{1}. Let σ_{1} = (a_{1} a_{2} ⋯ a_{n}).

Let T be the set S \ {a_{1}, a_{2}, …, a_{n}} consisting of all elements of S other than a_{1}, a_{2}, …, a_{n}. Now a_{1} = p(a_{n}), and a_{i} = p(a_{i−1}) for i = 2, 3, …, n. Thus if x ∈ T then p(x) ≠ a_{i} for i = 1, 2, …, n (since the function p: S → S is injective), and therefore p(x) ∈ T. We can therefore define a function q: T → T, where q(x) = p(x) for all x ∈ T. This function has a welldefined inverse q^{−1}: T → T where q^{−1}(x) = p^{−1}(x) for all x ∈ T. It follows that q: T → T is a permutation of T. The induction hypothesis ensures that this permutation is the identity permutation of T, or is a cycle, or can be expressed as a composition of two or more disjoint cyles. These cycles extend to permutations of S that fix the elements a_{1}, a_{2}, …, a_{n}, and these permutations of S are also cycles. It follows that either p = σ_{1} (and q is the identity permutation of T), or else p = σ_{1}σ_{2} … σ_{m}, where σ_{2}, σ_{3}, …, σ_{m} are disjoint cycles of S that fix a_{1}, a_{2}, …, a_{n} and correspond to cycles of T. Thus if the result holds for permutations of sets with fewer than k elements, then it holds for permutations of sets with k elements. It follows by induction on k that the result holds for permutations of finite sets.

Recall that a transposition is a permutation (a b) of a set S that interchanges two elements a and b of S and fixes the remaining elements.

Lemma 2.29 Every permutation of a finite set with more than one element can be expressed as a finite composition of transpositions.

Proof Each cycle can be expressed as a composition of transpositions. Indeed if a_{1}, a_{2}, …, a_{n} are distinct elements of a finite set S then

(a_{1} a_{2} ⋯ a_{n}) = (a_{1} a_{2})(a_{2} a_{3}) ⋯ (a_{n−1} a_{n}).

It follows from Proposition 2.28 that a permutation of S that is not the identity permutation can be expressed as a finite composition of transpositions. Moreover the identity permutation of S can be expressed as the composition of any transposition with itself, provided that S has more than one element. The result follows.

Theorem 2.30 A permutation of a finite set cannot be expressed in one way as a composition of an odd number of transpositions and in another way as a composition of an even number of transpositions.

Proof We can identify the finite set with the set {1, 2, …, n}, where n is the number of elements in the finite set. Let F: ℤ^{n} → ℤ be the function sending each n-tuple (m_{1}, m_{2}, …, m_{n}) of integers to the product ∏_{1≤j<k≤n}(m_{k} −m_{j}) of

the quantities m_{k} − m_{j} for all pairs (j, k) of integers satisfying 1 ≤ j < k ≤ n. Note that F (m_{1}, m_{2}, …, m_{n}) ≠ 0 whenever the integers m_{1}, m_{2}, …, m_{n} are distinct. If we transpose two of the integers m_{1}, m_{2}, …, m_{n} then this changes the sign of the function F, since the number of factors of the product ∏_{1≤j<k≤n}(m_{k} − m_{j}) that change sign is odd. (Indeed if we transpose m_{s} and

m_{t}, where 1 ≤ s < t < n then the factor m_{t} − m_{s} changes sign, the factor m_{t} − m_{i} becomes −(m_{i} − m_{s}) and the factor m_{i} − m_{s} becomes −(m_{t} − m_{i}) for each integer i for which s < i < t.) But any permutation σ of the set {1, 2, …, n} is a composition of transpositions. It follows that to each permutation σ of {1, 2, …, n} there corresponds a number ϵ_{σ}, where ϵ_{σ} = +1 or −1, such that F (m_{σ(1)}, m_{σ(2)}, …, m_{σ(n)}) = ϵ_{σ}F (m_{1}, m_{2}, …, m_{n}) for all integers m_{1}, m_{2}, …, m_{n}. Moreover ϵ_{στ} = ϵ_{σ}ϵ_{τ} for all permutations σ and τ of the set {1, 2, …, n}. Also ϵ_{τ} = −1 if the permutation τ is a transposition. It follows that if σ is expressible as a composition of r transpositions then ϵ_{σ} = (−1)^{r}. If σ is also expressible as a composition of s transpositions then ϵ_{σ} = (−1)^{s}, and hence (−1)^{r} = (−1)^{s}. But then r − s must be divisible by 2. The result follows.

A permutation of a finite set is said to be even if it is expressible as the composition of an even number of transpositions. A permutation of a finite set is said to be odd if it is expressible as the composition of an odd number of transpositions.

Any permutation of a finite set is expressible as a composition of transpositions (Lemma 2.29) and must therefore be either even or odd. However Theorem 2.30 ensures that a permutation of a finite set cannot be both even and odd.

Lemma 2.31 An n-cycle is even if n is odd, and is odd if n is even.

Proof An n-cycle (a_{1}, a_{2}, …, a_{n}) is expressible as a composition of n − 1 transpositions, since

(a_{1} a_{2} ⋯ a_{n}) = (a_{1} a_{2})(a_{2} a_{3}) ⋯ (a_{n−1} a_{n}).

Thus an n-cycle is even if n − 1 is even, and is odd if n − 1 is odd.

Example Let us classify the permutations of a set {a, b, c, d} of 4 elements into even and odd permutations. The identity permutation is even. The six transpositions are all odd. The eight 3-cycles are all even. The six 4cycles are all odd. The three remaining permutations (a b)(c d), (a c)(b d) and (a d)(b c) are all even. Note that there are 12 even permutations and 12 odd permutations of a set with 4 elements.

2.15 The Alternating Groups

A permutation of a finite set X is said to be even if it is the product of an even number of transpositions; it is said to be odd if it is the product of an odd number of transpositions. Note that the inverse of an even permutation and all products of even permutations are themselves even permutations.

Definition For each integer n satisfying n > 1, the alternating group A_{n} is the subgroup of the symmetric group Σ_{n} consisting of all even permutations of the set {1, 2, …, n}.

Note that, for each integer n satisfying n > 1, the alternating group A_{n} is a normal subgroup of Σ_{n} of index 2.

Example The alternating group A_{3} consists of the identity permutation and the cycles (1 2 3) and (1 3 2), and is thus isomorphic to the cyclic group C_{3} of order 3.

Lemma 2.32 Every even permutation of a finite set with at least three elements can be expressed as a product of cycles of order 3.

Proof Let X be a finite set. Then (a b)(b c) = (a b c) and (a b)(c d) = (c a d)(a b c) for all distinct elements a, b, c and d of X. Therefore the product of any two transpositions can be expressed as a product of cycles of order 3. The result thus follows from the fact that an even permutation is the product of an even number of transpositions.

Lemma 2.33 All cycles of order k in the alternating group A_{n} are conjugate to one another, provided that k ≤ n − 2.

Proof Let (m_{1} m_{2} ⋯ m_{k}) be a cycle of order k in A_{n}. Then there exists a permutation ρ of {1, 2, …, n} with the property that ρ(i) = m_{i} for i = 1, 2, …, k. If k ≤ n − 2 then any odd permutation with this property can be composed with the transposition that interchanges n − 1 and n to obtain an even permutation ρ with the required property. Then (m_{1} m_{2} ⋯ m_{k}) = ρ(1 2 ⋯ k)ρ^{−1}. Thus if k ≤ n−2 then all cycles of order k in A_{n} are conjugate to (1 2 ⋯ k) and are therefore conjugate to one another, as required.

Example We find all normal subgroups of the alternating group A_{4}. Let

V_{4} = {ι, (1 2)(3 4), (1 3)(2 4), (1 4)(2 3)},

where ι is the identity permutation. If ρ is an even permutation sending i to m_{i} for i = 1, 2, 3, 4 then ρ(1 2)(3 4)ρ^{−1} = (m_{1} m_{2})(m_{3} m_{4}). Therefore the permutations (1 2)(3 4), (1 3)(2 4) and (1 4)(2 3) are conjugate to one another, and hence V_{4} is a normal subgroup of A_{4} of order 4. The group V_{4} is referred to as the Kleinsche Viergruppe. It is isomorphic to the direct product C_{2}×C_{2} of two cyclic groups of order 2.

Let N be a normal subgroup of A_{4} that contains a cycle (m_{1} m_{2} m_{3}) of order 3. Now ρ(m_{1} m_{2} m_{3})ρ^{−1} = (ρ(m_{1}) ρ(m_{2}) ρ(m_{3})) for all ρ ∈ A_{4}. Therefore any cycle in A_{4} of order 3 is conjugate to either (m_{1} m_{2} m_{3}) or to (m_{1} m_{3} m_{2}). But (m_{1} m_{3} m_{2}) ∈ N, since (m_{1} m_{3} m_{2}) = (m_{1} m_{2} m_{3})^{2}. Therefore every cycle of order 3 in A_{4} belongs to the normal subgroup N. But then N = A_{4}, since A_{4} is generated by cycles of order 3 (Lemma 2.32). We have thus shown that if a normal subgroup N of A_{4} contains a cycle of order 3 then N = A_{4}.

Now let N be a normal subgroup of A_{4} that does not contain any cycle of order 3. Then N ⊂ V_{4}, since all elements of A_{4} \ V_{4} are cycles of order 3. But the only normal subgroups of A_{4} that are contained in V_{4} are {ι} and V_{4} itself, since the three elements of V_{4} \ {ι} are conjugate to one another.

We conclude that the normal subgroups of A_{4} are the trivial group {ι}, the Kleinsche Viergruppe V_{4} and A_{4} itself.

We recall that a non-trivial group G is simple if and only if the only normal subgroups of G are G itself and the trivial subgroup whose only element is the identity element of G. The alternating group A_{4} is not simple. We shall prove that A_{n} is simple when n ≥ 5.

Lemma 2.34 Let N be a non-trivial normal subgroup of the alternating group A_{n}, where n ≥ 5. Then there exists σ ∈ N, where σ is not the identity permutation, and a ∈ {1, 2, …, n} such that σ(a) = a.

Proof Let X = {1, 2, …, n}. The proof divides into two cases, depending on whether or not the normal subgroup N contains a permutation ρ of X with the property that ρ^{2} is not the identity permutation.

Suppose that the normal subgroup N contains a permutation ρ of X with the property that ρ^{2} is not the identity permutation. Then there exists a ∈ X such that ρ(ρ(a)) ≠ a. Let b = ρ(a) and c = ρ(b). Then the elements a, b and c are distinct. Choose elements d and e of X such that a, b, c, d and e are distinct. (This is possible since the set X has n elements, where n ≥ 5.) Let ρ′ = (c d e)ρ(c d e)^{−1}. Then ρ′ ∈ N (since ρ ∈ N and N is a normal subgroup), ρ′(a) = b and ρ′(b) = d. Now ρ′ ≠ ρ, since ρ′(b) ≠ ρ(b). Thus if σ = ρ^{−1}ρ′ then σ ∈ N, σ(a) = a, and σ is not the identity permutation.

It remains to prove the result in the case where ρ^{2} is the identity permutation for all ρ ∈ N. In this case choose ρ ∈ N, where ρ is not the identity permutation, let a be an element of X for which ρ(a) ≠ a, and let b = ρ(a). The permutation ρ is even (since it belongs to the alternating group A_{n}), and therefore ρ cannot be the transposition (a b). It follows that there exists an element c, distinct from a and b, such that ρ(c) ≠ c. Let d = ρ(c). Then the elements a, b, c and d of X are distinct. Choose an element e of X which is distinct from a, b, c and d. (This is possible since the set X has n elements, where n ≥ 5.) Let ρ′ = (c d e)ρ(c d e)^{−1}. Then ρ′(a) = b and ρ′(d) = e. Now ρ′ ≠ ρ, since ρ′(d) ≠ ρ(d). Thus if σ = ρ^{−1}ρ′ then σ ∈ N, σ(a) = a, and σ is not the identity permutation.

Lemma 2.35 Let N be a normal subgroup of A_{n}, where n ≥ 5. If N contains a 3-cycle then N = A_{n}.

Proof Suppose that N contains a 3-cycle. Then N contains every 3-cycle of A_{n}, since all 3-cycles in A_{n} are conjugate (Lemma 2.33). But then N contains every even permutation, since every even permutation is the identity permutation, a 3-cycle or a finite product of 3-cycles (Lemma 2.32). Thus N = A_{n}.

Theorem 2.36 The alternating group A_{n} is simple when n ≥ 5.

Proof First we prove that A_{5} is simple. Let N be a non-trivial normal subgroup of A_{5}. We shall show that N ∩ H is non-trivial, where H = {ρ ∈ A_{5} : ρ(5) = 5}. Then H is a subgroup of A_{5} that is isomorphic to A_{4}.

It follows from Lemma 2.34 that there exists σ ∈ N, where σ is not the identity permutation, and a ∈ {1, 2, 3, 4, 5} such that σ(a) = a. Choose ρ ∈ A_{5} such that ρ(a) = 5, and let σ′ = ρσρ^{−1}. Then σ′ ∈ N and σ′(5) = 5, and therefore σ′ ∈ H ∩ N. But σ′ is not the identity permutation. Thus H ∩ N is a non-trivial normal subgroup of H. But the subgroup H of A_{5} is isomorphic to A_{4} (since each permutation of {1, 2, 3, 4} can be regarded as a permutation of {1, 2, 3, 4, 5} that fixes 5). It follows from this that H ∩ N must contain the permutations (1 2)(3 4), (1 3)(2 4) and (1 4)(2 3) since the two non-trivial normal subgroups of A_{4} each contain these permutations. But then the normal subgroup N of A_{5} contains also the permutation (1 2)(4 5), since (1 2)(4 5) = (3 4 5)(1 2)(3 4)(3 4 5)^{−1}. It follows that N contains the cycle (3 4 5), since (3 4 5) = (1 2)(3 4)(1 2)(4 5). It follows from Lemma 2.35 that N = A_{5}. Thus the group A_{5} is simple.

We now prove that A_{n} is simple for n > 5 by induction on n. Thus suppose that n > 5 and the group A_{n−1} is simple. Let N be a non-trivial normal subgroup of A_{n}, and let H = {ρ ∈ A_{n} : ρ(n) = n}. It follows from Lemma 2.34 that there exists σ ∈ N, where σ is not the identity permutation, and a ∈ {1, 2, …, n} such that σ(a) = a. Choose ρ ∈ A_{n} such that ρ(a) = n, and let σ′ = ρσρ^{−1}. Then σ′ ∈ N and σ′(n) = n, and therefore σ′ ∈ H ∩ N. But σ′ is not the identity permutation. Thus H ∩ N is a non-trivial normal subgroup of H. But the subgroup H of A_{n} is simple, since it is isomorphic to A_{n−1}. It follows that N ∩ H = H, and thus H ⊂ N. But then N contains a 3-cycle, and therefore N = A_{n} (Lemma 2.35). Thus the group A_{n} is simple. We conclude by induction on n that the group A_{n} is simple whenever n ≥ 5, as required.

2.16 Normal Subgroups of the Symmetric Groups

We can now find all normal subgroups of the symmetric groups Σ_{n}. If N is a normal subgroup of Σ_{n} then N ∩ A_{n} is a normal subgroup of A_{n}. Moreover it follows from the First Isomorphism Theorem (Theorem 2.23) that N/(N ∩ A_{n}) ≅ NA_{n}/A_{n}. But NA_{n}/A_{n} is a subgroup of Σ_{n}/A_{n}, and |Σ_{n}/A_{n}| = 2. Therefore either N ⊂ A_{n} or else N ∩ A_{n} is a subgroup of N of index 2.

Example We show that if n ≥ 5 then the only normal subgroups of Σ_{n} are the trivial subgroup, the alternating group A_{n} and Σ_{n} itself. Now these subgroups are all normal subgroups of Σ_{n}. Moreover the trivial subgroup and A_{n} are the only normal subgroups of Σ_{n} contained in A_{n}, since A_{n} is simple when n ≥ 5 (Theorem 2.36).

Let N be a normal subgroup of Σ_{n} that is not contained in A_{n}. Then N ∩ A_{n} is a normal subgroup of A_{n}. Now if N ∩ A_{n} were the trivial subgroup then N would be a subgroup of Σ_{n} of order 2. But one can readily verify that Σ_{n} contains no normal subgroup of order 2 unless n = 2, in which case A_{2} is itself the trivial group. It follows that N ∩ A_{n} = A_{n}, and hence N = Σ_{n}. We have therefore shown that if n ≥ 5 then the only normal subgroups of Σ_{n} are the trivial subgroup, the alternating group A_{n} and Σ_{n} itself.

Example We now show that the only normal subgroups of the symmetric group Σ_{4} are the trivial subgroup, the Kleinsche Viergruppe V_{4}, the alternating group A_{4} and Σ_{4} itself.

The trivial group and the groups V_{4} and A_{4} are normal subgroups of Σ_{4}. Moreover they are the only normal subgroups of Σ_{4} contained in A_{4}, since they are the only normal subgroups of A_{4}.

Let N be a normal subgroup of Σ_{4} that is not contained in A_{4}. Then N ∩ A_{4} is a normal subgroup of A_{4}. One can readily verify that Σ_{4} contains no normal subgroup of order 2. It follows that V_{4} ⊂ N, since only normal subgroups of A_{4} other than the trivial subgroup are the groups V_{4} and A_{4}. Now the only odd permutations in Σ_{4} are transpositions and cycles of order 4.

Moreover if N contains a cycle of order 4 then N contains a transposition, since V_{4} ⊂ N and

(m_{1} m_{2})(m_{3} m_{4})(m_{1} m_{2} m_{3} m_{4}) = (m_{2} m_{4})

for all cycles (m_{1} m_{2} m_{3} m_{4}) of order 4. It follows that if N is a normal subgroup of Σ_{4} that is not contained in A_{4} then N must contain at least one transposition. But then N contains all transpositions, and therefore N = Σ_{4}. This shows that the only normal subgroups of Σ_{4} are the trivial group, the Kleinsche Viergruppe V_{4}, the alternating group A_{4} and Σ_{4} itself.

2.17 Finitely Generated Abelian Groups

Let H be a subgroup of the additive group ℤ^{n} consisting of all n-tuples of integers, with the operation of (vector) addition. A list b_{1}, b_{2}, …, b_{r} of elements of ℤ^{n} is said to constitute an integral basis (or ℤ-basis) of H if the following conditions are satisfied:

• the element m_{1}b_{1} + m_{2}b_{2} + ⋯ + m_{r}b_{r} belongs to H for all integers m_{1}, m_{2}, …, m_{r};

• given any element h of H, there exist uniquely determined integers m_{1}, m_{2}, …, m_{r} such that h = m_{1}b_{1} + m_{2}b_{2} + ⋯ + m_{r}b_{r}.

Note that elements b_{1}, b_{2}, …, b_{n} of ℤ^{n} constitute an integral basis of ℤ^{n} if and only if every element of ℤ^{n} is uniquely expressible as a linear combination of b_{1}, b_{2}, …, b_{n} with integer coefficients. It follows from basic linear algebra that the rows of an n × n matrix of integers constitute an integral basis of ℤ^{n} if and only if the determinant of that matrix is ±1.

Theorem 2.37 Let H be a non-trivial subgroup of ℤ^{n}. Then there exists an integral basis b_{1}, b_{2}, …, b_{n} of ℤ^{n}, a positive integer s, where s ≤ n, and positive integers k_{1}, k_{2}, …, k_{s} for which k_{1}b_{1}, k_{2}b_{2}, …, k_{s}b_{s} is an integral basis of H.

Proof We prove the result by induction on n. The result is clearly true when n = 1, since every non-trivial subgroup of ℤ is of the form kℤ for some positive integer k. Suppose therefore that n > 1 and that the result holds for all subgroups of ℤ^{n−1}. We must show that the result then holds for all subgroups H of ℤ^{n}.

Let k_{1} be the smallest strictly positive integer for which there exists some integral basis u_{1}, u_{2}, …, u_{n} of ℤ^{n} and some element of H of the form m_{1}u_{1} + m_{2}u_{2} + ⋯ + m_{n}u_{n} where m_{1}, m_{2}, …, m_{n} are integers and m_{i} = k_{1} for some integer i satisfying 1 ≤ i ≤ n. Let u_{1}, u_{2}, …, u_{n} be such a basis, with i = 1, and let h_{0} be an element of H for which h_{0} = m_{1}u_{1} + m_{2}u_{2} + ⋯ + m_{n}u_{n}, where m_{1}, m_{2}, …, m_{n} are integers and m_{1} = k_{1}.

We show that each coefficient m_{i} is divisible by k_{1}. Now, for each i, there exist integers q_{i} and r_{i} such that m_{i} = q_{i}k_{1} + r_{i} and 0 ≤ r_{i} < k_{1}. Let b_{1} = u_{1} + ∑_{i=2}^{n} q_{i}u_{i}. Then b_{1}, u_{2}, …, u_{n} is an integral basis of ℤ^{n} and

h_{0} = k_{1}b_{1} + ∑_{i=2}^{n} r_{i}u_{i}.

The choice of k_{1} now ensures that the coefficients r_{i} cannot be strictly positive (as they are less than k_{1}), and therefore r_{i} = 0 and m_{i} = q_{i}k_{1} for i = 2, 3, …, n. Moreover h_{0} = k_{1}b_{1}.

Now let φ: ℤ^{n−1} → ℤ^{n} be the injective homomorphism sending each element (m_{2}, m_{3}, …, m_{n}) of ℤ^{n−1} to ∑_{i=2}^{n} m_{i}u_{i}, and let H̃ = φ^{−1}(H). Then, given any element h of H, there exist an integer m and an element h̃ of ℤ^{n−1} such that h = mb_{1} + φ(h̃). Moreover m and h̃ are uniquely determined by h, since b_{1}, u_{2}, …, u_{n} is an integral basis of ℤ^{n}. Let m = qk_{1} + r, where q and r are integers and 0 ≤ r < k_{1}. Then h − qh_{0} = rb_{1} + φ(h̃), where φ(h̃) is expressible as a linear combination of u_{2}, …, u_{n} with integer coefficients. The choice of k_{1} now ensures that r cannot be strictly positive, and therefore r = 0. Then φ(h̃) ∈ H, and hence h̃ ∈ H̃. We conclude from this that, given any element h of H, there exist an integer q and an element h̃ of H̃ such that h = qk_{1}b_{1} + φ(h̃). Moreover q and h̃ are uniquely determined by h.

Now the induction hypothesis ensures the existence of an integral basis b̃_{2}, b̃_{3}, …, b̃_{n} of ℤ^{n−1} for which there exist positive integers k_{2}, k_{3}, …, k_{s} such that k_{2}b̃_{2}, k_{3}b̃_{3}, …, k_{s}b̃_{s} is an integral basis of H̃. Let b_{i} = φ(b̃_{i}) for each integer i between 2 and n. One can then readily verify that b_{1}, b_{2}, …, b_{n} is an integral basis of ℤ^{n} and k_{1}b_{1}, k_{2}b_{2}, …, k_{s}b_{s} is an integral basis of H, as required.

An Abelian group G is generated by elements g_{1}, g_{2}, …, g_{n} if and only if every element of G is expressible in the form g_{1}^{m_{1}}g_{2}^{m_{2}} ⋯ g_{n}^{m_{n}} for some integers m_{1}, m_{2}, …, m_{n}.

Lemma 2.38 A non-trivial Abelian group G is finitely generated if and only if there exists a positive integer n and some surjective homomorphism θ: ℤ^{n} → G.

Proof Let e_{1}, e_{2}, …, e_{n} be the integral basis of ℤ^{n} with e_{1} = (1, 0, …, 0), e_{2} = (0, 1, 0, …, 0), …, e_{n} = (0, …, 0, 1). If there exists a surjective homomorphism θ: ℤ^{n} → G then G is generated by g_{1}, g_{2}, …, g_{n}, where g_{i} = θ(e_{i}) for i = 1, 2, …, n. Conversely if G is generated by g_{1}, g_{2}, …, g_{n} then there is a surjective homomorphism θ: ℤ^{n} → G that sends (m_{1}, m_{2}, …, m_{n}) ∈ ℤ^{n} to g_{1}^{m_{1}}g_{2}^{m_{2}} ⋯ g_{n}^{m_{n}}.

Theorem 2.39 Let G be a non-trivial finitely generated Abelian group. Then there exist a positive integer n and a non-negative integer s between 0 and n, such that if s = 0 then G ≅ ℤ^{n}, and if s > 0 then there exist positive integers k_{1}, k_{2}, …, k_{s} such that

G ≅ C_{k_{1}} × C_{k_{2}} × ⋯ × C_{k_{s}} × ℤ^{n−s},

where C_{k_{i}} is a cyclic group of order k_{i} for i = 1, 2, …, s.

Proof There exists a positive integer n and some surjective homomorphism θ: ℤ^{n} → G, since G is finitely-generated. Let H be the kernel of θ. If H is trivial then the homomorphism θ is an isomorphism between ℤ^{n} and G. If H is non-trivial then G is isomorphic to ℤ^{n}/H, and there exists an integral basis b_{1}, b_{2}, …, b_{n} of ℤ^{n}, a positive integer s, where s ≤ n, and positive integers k_{1}, k_{2}, …, k_{s} for which k_{1}b_{1}, k_{2}b_{2}, …, k_{s}b_{s} is an integral basis of H (Theorem 2.37). Then the group ℤ^{n}/H, and thus G, is isomorphic to C_{k_{1}} × C_{k_{2}} × ⋯ × C_{k_{s}} × ℤ^{n−s}, where C_{i} is a cyclic group of order k_{i} for i = 1, 2, …, s. Indeed there is a well-defined homomorphism φ: ℤ^{n} → C_{k_{1}} × C_{k_{2}} × ⋯ × C_{k_{s}} × ℤ^{n−s} which sends each element

m_{1}b_{1} + m_{2}b_{2} + ⋯ + m_{n}b_{n}

of ℤ^{n} to (a_{1}^{m_{1}}, a_{2}^{m_{2}}, …, a_{s}^{m_{s}}, m_{s+1}, …, m_{n}), where a_{i} is a generator of the cyclic group C_{i} for i = 1, 2, …, s. The homomorphism φ is surjective, and its kernel is the subgroup H. Therefore G ≅ ℤ^{n}/H ≅ C_{k_{1}}×C_{k_{2}}×⋯×C_{k_{s}}×ℤ^{n−s}, as required.

Corollary 2.40 Let G be a non-trivial finite Abelian group. Then there exist positive integers k_{1}, k_{2}, …, k_{n} such that G ≅ C_{k_{1}} ×C_{k_{2}} ×⋯×C_{k_{n}}, where C_{k_{i}} is a cyclic group of order k_{i} for i = 1, 2, …, n.

With some more work it is possible to show that the positive integers k_{1}, k_{2}, …, k_{s} in Theorem 2.39 may be chosen such that k_{1} > 1 and k_{i−1} divides k_{i} for i = 2, 3, …, s, and that the Abelian group is then determined up to isomorphism by the integer n and the sequence of positive integers k_{1}, k_{2}, …, k_{s}.

2.18 The Class Equation of a Finite Group

Definition The centre Z(G) of a group G is the subgroup of G defined by

Z(G) = {g ∈ G : gh = hg for all h ∈ G}.

One can verify that the centre of a group G is a normal subgroup of G. Let G be a finite group, and let Z(G) be the centre of G. Then G \ Z(G)

is a disjoint union of conjugacy classes. Let r be the number of conjugacy classes contained in G\Z(G), and let n_{1}, n_{2}, …, n_{r} be the number of elements in these conjugacy classes. Then n_{i} > 1 for all i, since the centre Z(G) of G is the subgroup of G consisting of those elements of G whose conjugacy class contains just one element. Now the group G is the disjoint union of its conjugacy classes, and therefore

|G| = |Z(G)| + n_{1} + n_{2} + ⋯ + n_{r}.

This equation is referred to as the class equation of the group G.

Definition Let g be an element of a group G. The centralizer C(g) of g is the subgroup of G defined by C(g) = {h ∈ G : hg = gh}.

Proposition 2.41 Let G be a finite group, and let p be a prime number. Suppose that p^{k} divides the order of G for some positive integer k. Then either p^{k} divides the order of some proper subgroup of G, or else p divides the order of the centre of G.

Proof Choose elements g_{1}, g_{2}, …, g_{r} of G\Z(G), where Z(G) is the centre of G, such that each conjugacy class included in G \ Z(G) contains exactly one of these elements. Let n_{i} be the number of elements in the conjugacy class of g_{i} and let C(g_{i}) be the centralizer of g_{i} for each i. Then C(g_{i}) is a proper subgroup of G, and |G| = n_{i}|C(g_{i})|. Thus if p^{k} divides |G| but does not divide the order of any proper subgroup of G then p must divide n_{i} for i = 1, 2, …, r. Examination of the class equation |G| = |Z(G)| + n_{1} + n_{2} + ⋯ + n_{r} now shows that p divides |Z(G)|, as required.

2.19 Cauchy’s Theorem

Theorem 2.42 (Cauchy) Let G be an finite group, and let p be a prime number that divides the order of G. Then G contains an element of order p.

Proof We prove the result by induction on the order of G. Thus suppose that every finite group whose order is divisible by p and less than |G| contains an element of order p. If p divides the order of some proper subgroup of G then that subgroup contains the required element of order p. If p does not divide the order of any proper subgroup of G then Proposition 2.41 ensures that p divides the order of the centre Z(G) of G, and thus Z(G) cannot be a proper subgroup of G. But then G = Z(G) and the group G is Abelian.

Thus let G be an Abelian group whose order is divisible by p, and let H be a proper subgroup of G that is not contained in any larger proper subgroup. If |H| is divisible by p then the induction hypothesis ensures that H contains the required element of order p, since |H| < |G|. Suppose then that |H| is not divisible by p. Choose g ∈ G \ H, and let C be the cyclic subgroup of G generated by g. Then HC = G, since HC ≠ H and HC is a subgroup of G containing H. It follows from the First Isomorphism Theorem (Theorem 2.23) that G/H ≅ C/H ∩ C. Now p divides |G/H|, since |G/H| = |G|/|H| and p divides |G| but not |H|. Therefore p divides |C|. Thus if m = |C|/p then g^{m} is the required element of order p. This completes the proof of Cauchy’s Theorem.

2.20 The Structure of p-Groups

Definition Let p be a prime number. A p-group is a finite group whose order is some power p^{k} of p.

Lemma 2.43 Let p be a prime number, and let G be a p-group. Then there exists a normal subgroup of G of order p that is contained in the centre of G.

Proof Let |G| = p^{k}. Then p^{k} divides the order of G but does not divide the order of any proper subgroup of G. It follows from Proposition 2.41 that p divides the order of the centre of G. It then follows from Cauchy’s Theorem (Theorem 2.42) that the centre of G contains some element of order p. This element generates a cyclic subgroup of order p, and this subgroup is normal since its elements commute with every element of G.

Proposition 2.44 Let G be a p-group, where p is some prime number, and let H be a proper subgroup of G. Then there exists some subgroup K of G such that H ◁ K and K/H is a cyclic group of order p.

Proof We prove the result by induction on the order of G. Thus suppose that the result holds for all p-groups whose order is less than that of G. Let Z be the centre of G. Then ZH is a well-defined subgroup of G, since Z is a normal subgroup of G.

Suppose that ZH ≠ H. Then H is a normal subgroup of ZH. The quotient group ZH/H is a p-group, and contains a subgroup K_{1} of order p (Lemma 2.43). Let K = {g ∈ ZH : gH ∈ K_{1}}. Then H ◁K and K/H ≅ K_{1}, and therefore K is the required subgroup of G.

Finally suppose that ZH = H. Then Z ⊂ H. Let H_{1} = {hZ : h ∈ H}. Then H_{1} is a subgroup of G/Z. But G/Z is a p-group, and |G/Z| < |G|, since |Z| ≥ p (Lemma 2.43). The induction hypothesis ensures the existence of a subgroup K_{1} of G/Z such that H_{1} ◁ K_{1} and K_{1}/H_{1} is cyclic of order p. Let K = {g ∈ G : gZ ∈ K_{1}}. Then H ◁ K and K/H ≅ K_{1}/H_{1}. Thus K is the required subgroup of G.

Repeated applications of Proposition 2.44 yield the following result.

Corollary 2.45 Let G be a finite group whose order is a power of some prime number p. Then there exist subgroups G_{0}, G_{1}, …, G_{n} of G, where G_{0} is the trivial subgroup and G_{n} = G, such that G_{i−1} ◁ G_{i} and G_{i}/G_{i−1} is a cyclic group of order p for i = 1, 2, …, n.

2.21 The Sylow Theorems

Definition Let G be a finite group, and let p be a prime number dividing the order |G| of G. A p-subgroup of G is a subgroup whose order is some power of p. A Sylow p-subgroup of G is a subgroup whose order is p^{k}, where k is the largest natural number for which p^{k} divides |G|.

Theorem 2.46 (First Sylow Theorem) Let G be a finite group, and let p be a prime number dividing the order of G. Then G contains a Sylow p-subgroup.

Proof We prove the result by induction on the order of G. Thus suppose that all groups whose order is less than that of G contain the required Sylow p-subgroups. Let k be the largest positive integer for which p^{k} divides |G|. If p^{k} divides the order of some proper subgroup H of G then the induction hypothesis ensures that H contains the required Sylow p-subgroup of order p^{k}. If p^{k} does not divide the order of any proper subgroup of G then p divides the order of the centre Z(G) of G (Proposition 2.41). It follows from Cauchy’s Theorem (Theorem 2.42) that Z(G) contains an element of order p, and this element generates a normal subgroup N of G of order p. The induction hypothesis then ensures that G/N has a Sylow p-subgroup L of order p^{k−1}, since |G/N| = |G|/p. Let K = {g ∈ G : gN ∈ L}. Then |K| = p|L| = p^{k}, and thus K is the required Sylow p-subgroup of G.

Theorem 2.47 (Second Sylow Theorem) Let G be a finite group, and let p be a prime number dividing the order of G. Then all Sylow p-subgroups of G are conjugate, and any p-subgroup of G is contained in some Sylow psubgroup of G. Moreover the number of Sylow p-subgroups in G divides the order of |G| and is congruent to 1 modulo p.

Proof Let K be a Sylow p-subgroup of G, and let X be the set of left cosets of K in G. Let H be a p-subgroup of G. Then H acts on X on the left, where h(gK) = hgK for all h ∈ H and g ∈ G. Moreover h(gK) = gK if and only if g^{−1}hg ∈ K. Thus an element gK of X is fixed by H if and only if g^{−1}Hg ⊂ K.

Let |G| = p^{k}m, where k and m are positive integers and m is coprime to p. Then |K| = p^{k}. Now the number of left cosets of K in G is |G|/|K|. Thus the set X has m elements. Now the number of elements in any orbit for the action of H on X divides the order of H, since it is the index in H of the stabilizer of some element of that orbit (Lemma 2.26). But then the number of elements in each orbit must be some power of p, since H is a p-group. Thus if an element of X is not fixed by H then the number of elements in its orbit is divisible by p. But X is a disjoint union of orbits under the action of H on X. Thus if m′ denotes the number of elements of X that are fixed by H then m − m′ is divisible by p.

Now m is not divisible by p. It follows that m′ ≠ 0, and m′ is not divisible by p. Thus there exists at least one element g of G such that g^{−1}Hg ⊂ K. But then H is contained in the Sylow p-subgroup gKg^{−1}. Thus every p-subgroup is contained in a Sylow p-subgroup of G, and this Sylow p-subgroup is a conjugate of the given Sylow p-subgroup K. In particular any two Sylow p-subgroups are conjugate.

It only remains to show that the number of Sylow p-subgroups in G divides the order of |G| and is congruent to 1 modulo p. On applying the above results with H = K, we see that g^{−1}Kg = K for some g ∈ G if and only if gK is a fixed point for the action of K on X. But the number of elements g of G for which gK is a fixed point is m′|K|, where m′ is the number of fixed points in X. It follows that the number of elements g of G for which g^{−1}Kg = K is p^{k}m′. But every Sylow p-subgroup of G is of the form g^{−1}Kg for some g ∈ G. It follows that the number n of Sylow p-subgroups in G is given by n = |G|/p^{k}m′ = m/m′. In particular n divides |G|. Now we have already shown that m − m′ is divisible by p. It follows that m′ is coprime to p, since m is coprime to p. Also m − m′ is divisible by m′, since (m − m′)/m′ = n − 1. Putting these results together, we see that m − m′ is divisible by m′p, and therefore n − 1 is divisible by p. Thus n divides |G| and is congruent to 1 modulo p, as required.

2.22 Some Applications of the Sylow Theorems

Theorem 2.48 Let p and q be prime numbers, where p < q and q ≢ 1 (mod p). Then any group of order pq is cyclic.

Proof Let G be a group of order pq. It follows from the First Sylow Theorem that G contains Sylow subgroups N_{p} and N_{q} of orders p and q respectively. Now the number n_{p} of Sylow p-subgroups divides pq and satisfies n_{p} ≡ 1 (mod p), by the Second Sylow Theorem. Clearly n_{p} cannot be divisible by p, and therefore either n_{p} = 1 or n_{p} = q. But q ≢ 1 (mod p). It follows that n_{p} = 1. Thus the group G has just one subgroup of order p.

Now, given any element g of G, the subgroups N_{p} and gN_{p}g^{−1} are of order p. It follows that gN_{p}g^{−1} = N_{p} for all elements g of G. Thus N_{p} is a normal subgroup of G.

A similar argument shows that N_{q} is also a normal subgroup of G, since p < q, and therefore p ≢ 1 (mod q).

Now N_{p} ∩N_{q} is a subgroup of both N_{p} and N_{q}. It follows from Lagrange’s Theorem that the order of N_{p} ∩ N_{q} divides both of the prime numbers p and q, and therefore |N_{p} ∩ N_{q}| = 1 and N_{p} ∩ N_{q} = {e}, where e is the identity element of G.

Let x ∈ N_{p} and y ∈ N_{q}. Then yx^{−1}y^{−1} ∈ N_{p} and xyx^{−1} ∈ N_{q}, since N_{p} and N_{q} are normal subgroups of G. But then xyx^{−1}y^{−1} ∈ N_{p} ∩ N_{q}, since xyx^{−1}y^{−1} = x(yx^{−1}y^{−1}) = (xyx^{−1})y^{−1}, and therefore xyx^{−1}y^{−1} = e. Thus xy = yx for all x ∈ N_{p} and y ∈ N_{q}. It follows easily from this that the function φ: N_{p} × N_{q} → G which sends (x, y) ∈ N_{p} × N_{q} to xy is a homomorphism. This homomorphism is injective, for if xy = e for some x ∈ N_{p} and y ∈ N_{q}, then x = y^{−1}, and hence x ∈ N_{p} ∩ N_{q}, from which it follows that x = e and y = e. But any injective homomorphism between two finite groups of the same order is necessarily an isomorphism. Therefore the function φ: N_{p} × N_{q} → G is an isomorphism, and thus G ≅ N_{p} × N_{q}.

Now any group whose order is prime number must be cyclic. Therefore the groups N_{p} and N_{q} are cyclic. Let x be an element of N_{p} that generates N_{p}, and let y be an element of N_{q} that generates N_{q}. Then (x, y)^{n} = (x^{n}, y^{n}) for all integers n. It follows from this that the order of (x, y) cannot be equal to 1, p or q, and must therefore be equal to pq. Thus N_{p} × N_{q} is a cyclic group generated by (x, y), and therefore G is a cyclic group, generated by xy, as required.

Example Any finite group whose order is 15, 33, 35, 51, 65, 69, 85, 87, 91 or 95 is cyclic.

Theorem 2.49 Let G be a group of order 2p where p is a prime number greater than 2. Then either the group G is cyclic, or else the group G is isomorphic to the dihedral group D_{2p} of symmetries of a regular p-sided polygon in the plane.

Proof It follows from the First Sylow Theorem, or from Cauchy’s Theorem, that the group G contains elements x and y whose orders are 2 and p respectively. The subgroup N generated by y is then a Sylow p-subgroup of G. Now it follows from the Second Sylow Theorem that the number of Sylow p-subgroups of G divides 2p and is congruent to 1 modulo p. There can therefore be only one such Sylow p-subgroup, since 2, p and 2p are not congruent to 1 modulo p. Now if g is any element of G then gNg^{−1} is a Sylow p-subgroup of G, and therefore gNg^{−1} = N. We deduce that N is a normal subgroup of G, of order p.

Now consider the element xyx^{−1} of G. This must be an element of the normal subgroup N of G generated by y. Therefore xyx^{−1} = y^{k} for some integer k. Moreover k is not divisible by p, since xyx^{−1} is not the identity element e of G. Then

y^{k^{2}} = (y^{k})^{k} = (xyx^{−1})^{k} = xy^{k}x^{−1} = x(xyx^{−1})x^{−1} = x^{2}yx^{−2}.

But x^{2} = x^{−2} = e, since x is an element of G of order 2. It follows that y^{k^{2}} = y, and thus y^{k^{2}−1} = e. But then p divides k^{2} − 1, since y is an element of order p. Moreover k^{2} − 1 = (k − 1)(k + 1). It follows that either p divides k − 1, in which case xyx^{−1} = y, or else p divides k + 1, in which case xyx^{−1} = y^{−1}.

In the case when xyx^{−1} = y we see that xy = yx, and one can then readily verify that the group G is a cyclic group of order 2p generated by xy.

In the case when xyx^{−1} = y^{−1} the group G is isomorphic to the dihedral group D_{2p} of order 2p. In this case the elements x and y generate G (since they generate a subgroup of G whose order divides 2p but is greater than p, and must therefore be equal to 2p). Under the isomorphism with the dihedral group D_{2p} the element x corresponds to a reflection in one of the axes of symmetry of the regular p-sided polygon, and the element y corresponds to a rotation of that polygon about its centre through an angle of 2π/p radians.

Theorem 2.50 Let p and q be prime numbers with p < q, and let d be the smallest positive integer for which p^{d} ≡ 1 (mod q). If G is a group of order p^{k}q, where 0 < k < d then G contains a normal subgroup of order q. If G is a group of order p^{d}q then either G contains a normal subgroup of order q or else G contains a normal subgroup of order p^{d}.

Proof It follows from the First Sylow Theorem (or directly from Cauchy’s Theorem) that the group G contains at least one Sylow q-subgroup K, and this is of order q. If |G| = p^{k}q then the number n_{q} of such Sylow q-subgroups divides p^{k}q and satisfies n_{q} ≡ 1 (mod q), by the Second Sylow Theorem. It follows that n_{q} is coprime to q, and therefore n_{q} = p^{j} for some integer j satisfying 0 ≤ j ≤ k.

If k < d then none of the integers p, p^{2}, …, p^{k} are congruent to 1 modulo q, and therefore j = 0 and n_{q} = 1. In this case there is just one Sylow q-subgroup K, and this is a normal subgroup. (Given any element g of G, the subgroup gKg^{−1} is a Sylow q-subgroup, and therefore gKg^{−1} = K.)

If k = d then none of the integers p^{j} with 0 < j < d are congruent to 1 modulo q, and therefore either n_{q} = 1 or n_{q} = p^{d}. If n_{q} = 1 then there is just one Sylow q-subgroup K, and this is a normal subgroup.

If n_{q} > 1 then n_{q} = p^{d}, and thus there are p^{d} Sylow q-subgroups, and these are of order q. Now if K_{i} and K_{j} are two distinct subgroups of order q then K_{i} ∩ K_{j} is a proper subgroup of both K_{i} and K_{j}, and its order is a proper divisor of the order q of K_{i} and K_{j}, by Lagrange’s Theorem. But q is a prime number. It follows that K_{i} ∩ K_{j} = {e}, where e is the identity element of G. We deduce from this that no element of G of order q can belong to more than one subgroup of order q. But each subgroup of G of order q contains q −1 elements of order q (namely all elements of that subgroup with the exception of the identity element). It follows that the group G contains p^{d}(q − 1) elements of order q. Now |G| = p^{d}q. It follows that G contains exactly p^{d} elements that are not of order q. But it follows from the First Sylow Theorem that G contains at least one Sylow p-subgroup H, and this is of order p^{d}. This subgroup must therefore contain all the elements of G that are not of order q. It follows that the group G cannot contain more than one such Sylow p-subgroup. This Sylow p-subgroup H is therefore a normal subgroup of G of order p^{d}, as required.

2.23 Simple Groups

Definition A non-trivial group G is said to be simple if the only normal subgroups of G are the whole of G and the trivial subgroup {e} whose only element is the identity element e of G.

Lemma 2.51 Any non-trivial Abelian simple group is a cyclic group whose order is a prime number.

Proof Let G be a non-trivial Abelian simple group, and let x be an element of G that is not equal to the identity element e of G. All subgroups of an Abelian group are normal subgroups. Therefore the subgroup of G generated by x is a normal subgroup of G, and must therefore be the whole of G. Therefore G is a cyclic group, generated by the element x. Moreover all elements of G other than the identity element are generators of G, and are therefore of order p, where p = |G|. Let d be a divisor of p. Then x^{d} is an element of order p/d, since p/d is the smallest positive integer k for which x^{dk} = e. It follows that either d = 1 or d = p (since the group G contains no element whose order is greater than 1 but less than p). It follows that the order p of G is a prime number, as required.

We have already proved that if n ≥ 5 then the alternating group A_{n} has no normal subgroups other than the trivial subgroup and the whole of A_{n}. Thus A_{n} is a simple group when n ≥ 5. The order of the group A_{n} is n!/2. In particular |A_{5}| = 60.

Using the Sylow Theorems and related results, we can prove that any finite simple group whose order is less than 60 is a cyclic group of prime order.

Now the prime numbers less than 60 are the following: 2, 3, 5, 7, 11, 13, 17, 19, 23, 29, 31, 37, 41, 43, 47, 53 and 59. All groups of these orders are simple groups, and are cyclic groups.

If p is a prime number greater than 2 then any group of order 2p is either a cyclic group or else is isomorphic to the dihedral group D_{2p} of order 2p (Theorem 2.49). In either case such a group contains a normal subgroup of order p, and therefore not a simple group. In particular, there are no simple groups of orders 6, 10, 14, 22, 26, 34, 38, 46 or 58.

If G is a group of order p^{k} for some prime number p and for some integer k satisfying k ≥ 2, then G contains a normal subgroup of order p (Lemma 2.43). It follows that such a group is not simple. In particular, there are no simple groups of orders 4, 8, 16, 32, 9, 27, 25 and 49.

Let G be a group of order pq, where p and q are prime numbers and p < q. Any Sylow q-subgroup of G is of order q, and the number of such Sylow q-subgroups must divide pq and be congruent to 1 modulo q. Now p cannot be congruent to 1 modulo q, since 1 < p < q. Therefore G has just one Sylow q-subgroup, and this is a normal subgroup of G of order q. It follows that such a group is not a simple group. In particular there are no simple groups of orders 15, 21, 33, 35, 39, 51, 55 or 57. (In particular it follows from Theorem 2.48 that any group whose order is 15, 33, 35, or 51 is a cyclic group.)

It only remains to verify that there are no simple groups of orders 12, 18, 20, 24, 28, 30, 36, 40, 42, 44, 45, 48, 50, 52, 54 or 56.

We can deal with many of these on applying Theorem 2.50. On applying this theorem with p = 2, q = 3 and d = 2, we see that there are no simple groups of orders 6 or 12. On applying the theorem with p = 2, q = 5 and d = 4, we see that there are no simple groups of orders 10, 20, 40 or 80. On applying the theorem with p = 2, q = 7 and d = 3, we see that there are no simple groups of orders 14, 28 or 56. On applying the theorem with p = 2, q = 11 we see that there are no simple groups of orders 22, 44 etc., on applying the theorem with p = 2, q = 13 we see that there are no simple groups of orders 26, 52 etc., and on applying the theorem with p = 3 and q = 5, we see that there are no simple groups of orders 15, 45 etc.

It now remains to verify that there are no simple groups of orders 18, 24, 30, 36, 42, 48, 50 or 54.

Using the Second Sylow Theorem, we see that any group of order 18 has just one Sylow 3-subgroup. This Sylow 3-subgroup is then a normal group of order 9, and therefore no group of order 18 is simple. Similarly a group of order 50 has just one Sylow 5-subgroup, which is then a normal subgroup of order 25, and therefore no group of order 50 is simple. Also a group of order 54 has just one Sylow 3-subgroup, which is then a normal subgroup of order 27, and therefore no group of order 54 is simple.

On applying the Second Sylow Theorem, we see the number of Sylow 7-subgroups of any group of order 42 must divide 42 and be congruent to 1 modulo 7. This number must then be coprime to 7 and therefore divide 6, since 42 = 7 × 6. But no divisor of 6 greater than 1 is coprime to 1 modulo 7. It follows that any group of order 42 has just one Sylow 7-subgroup, and this subgroup is therefore a normal subgroup of order 7. Thus no group of order 42 is simple.

On applying the Second Sylow Theorem, we see that if a group of order 30 has more than one subgroup of order 3 then it must have 10 such subgroups, and must therefore have 20 elements of order 3 (since each subgroup of order 3 contains two elements of order 3, and the intersection of two distinct subgroups of order 3 must be the trivial subgroup). Similarly if a group of order 30 has more than one subgroup of order 5 then it must have 6 such subgroups, and must therefore have 24 elements of order 5. Obviously such a group cannot have both 20 elements of order 3 and 24 elements of order 5. Therefore it either has a single subgroup of order 3 or a single subgroup of order 5. This subgroup is normal. Therefore no group of order 30 is simple.

In order to show that there are no simple groups of order less than 60, apart from the cyclic groups whose order is prime, it only remains to verify that there are no simple groups of orders 24, 36 and 48. In order to deal with these remaining cases, we need to make use of the following result.

Lemma 2.52 Let H and K be subgroups of a finite group G. Then

|H ∩ K| ≥ \frac{|H| |K|}{|G|}.

Proof Let φ: H × K → G be the function with φ(h, k) = hk for all h ∈ H and k ∈ K. (This function is not in general a homomorphism.) Let (h_{1}, k_{1}) and (h_{2}, k_{2}) be elements of H × K. Then h_{1}k_{1} = h_{2}k_{2} if and only if h_{2}^{−1}h_{1} = k_{2}k_{1}^{−1}, in which case h_{2}^{−1}h_{1} ∈ H ∩ K. But then h_{2} = h_{1}x^{−1} and k_{2} = xk_{1} for some element x of H ∩ K. Thus φ(h_{1}, k_{1}) = φ(h_{2}, k_{2}) if and only if (h_{2}, k_{2}) = (h_{1}x^{−1}, xk_{1}) for some element x of H ∩ K. It follows that each element of the range φ(H × K) of the function φ is the image of exactly

|H ∩ K| elements of H × K. It follows from this that φ(H × K) has \frac{|H| |K|}{|H ∩ K|}

elements. But φ(H × K) is a subset of G. Therefore

\frac{|H| |K|}{|H ∩ K|} ≤ |G|.

The required inequality now follows directly.

Let G be a finite group, and let H be a subgroup of index 2 in G (i.e., a subgroup with half as many elements as G). Then H is a normal subgroup of G. Indeed the subsets H and G \ H of G are the left cosets and are also the right cosets of H in G, and therefore the left cosets of H in G coincide with the right cosets.

Example We now show that there are no simple groups of order 24. Let G be a group of order 24. Then G contains a Sylow 2-subgroup H of order 8. If this is the only Sylow 2-subgroup, then it is a normal subgroup, and therefore the group G is not simple. Otherwise the group G contains at least two distinct subgroups H and K of order 8. It then follows from Lemma 2.52 that |H ∩ K| ≥ \frac{8}{3}. But |H ∩ K| divides 8, by Lagrange’s Theorem, since H ∩ K is a subgroup of H and of K. Therefore |H ∩ K| = 4. It follows that H ∩ K is a subgroup of index 2 in H and K, and is therefore a normal subgroup of both H and K. Let

J = {g ∈ G : g(H ∩ K)g^{−1} = H ∩ K}.

Then J is a subgroup of G, and H ∩ K is a normal subgroup of J. Moreover H and K are subgroups of J, and therefore |J| is divisible by 8, by Lagrange’s Theorem. But J is a subgroup of G, and hence |J| divides 24. Also |J| > 8, since H (and K) are proper subgroups of J. It follows that |J| = 24, and therefore J = G. But then H ∩ K is a normal subgroup of G of order 4, and therefore G is not simple.

An analogous argument shows that there are no simple groups of order 48: a group G of order 48 contains either a single Sylow 2-subgroup of order 16, which is then a normal subgroup of G, or else it contains a normal subgroup of order 8 which is the intersection of two distinct Sylow 2-subgroups of G.

The following result will be needed in order to show that there are no simple groups of order 36. (It may be obtained as an immediate corollary of Proposition 2.44.)

Lemma 2.53 Let G be a group of order p^{2} where p is a prime number, and let H be a subgroup of G of order p. Then H is a normal subgroup of G.

Proof Let J = {g ∈ G : gHg^{−1} = H}. Then J is a subgroup of G and H is a normal subgroup of J. We shall show that J = G.

Now the centre Z(G) of G is contained in J. Moreover it follows from Lemma 2.43 that |Z(G)| is divisible by p. Were it the case that |J| = p then J = H = Z(G). But then J would consist of all elements of G for which gZ(G)g^{−1} = Z(G), and thus would be the whole of G, which is impossible. It follows that |J| = p^{2} (since |J| > p and |J| divides p^{2}). But then J = G, and hence H is a normal subgroup of G, as required.

Example We now show that there are no simple groups of order 36. Let G be a group of order 36. Then G contains a Sylow 3-subgroup H of order 9. If this is the only Sylow 3-subgroup, then it is a normal subgroup, and therefore the group G is not simple. Otherwise the group G contains at least two distinct subgroups H and K of order 9. It then follows from Lemma 2.52 that |H ∩ K| ≥ \frac{9}{4}. But |H ∩ K| divides 9, by Lagrange’s Theorem, since H ∩ K is a subgroup of H and of K. Therefore |H ∩ K| = 3. On applying Lemma 2.53 we see that H ∩ K is a normal subgroup of H and of K.

Let

J = {g ∈ G : g(H ∩ K)g^{−1} = H ∩ K}.

Then J is a subgroup of G, and H ∩ K is a normal subgroup of J. Moreover H and K are subgroups of J, and therefore |J| is divisible by 9, by Lagrange’s Theorem. But J is a subgroup of G, and hence |J| divides 36. Also |J| > 9, since H (and K) are proper groups of J. It follows that either |J| = 18 or 36. If |J| = 36 then J = G and H ∩ K is a normal subgroup of G of order 3. If |J| = 18 then J is a subgroup of G of index 2, and is therefore a normal subgroup of order 18. We conclude that any group of order 36 contains at least one non-trivial normal subgroup. Therefore there are no simple groups of order 36.

We have now shown that there are indeed no simple groups of order less than 60, other than the cyclic groups of prime order.

2.24 Solvable Groups

Definition A group G is said to be solvable (or soluble) if there exists a finite sequence G_{0}, G_{1}, …, G_{n} of subgroups of G, where G_{0} = {1} and G_{n} = G, such that G_{i−1} is normal in G_{i} and G_{i}/G_{i−1} is Abelian for i = 1, 2, …, n.

Example The symmetric group Σ_{4} is solvable. Indeed let V_{4} be the Kleinsche Viergruppe consisting of the identity permutation ι and the permutations (12)(34), (13)(24) and (14)(23), and let A_{4} be the alternating group consisting of all even permutations of {1, 2, 3, 4}. Then {ι} ◁ V_{4} ◁ A_{4} ◁ Σ_{4}, V_{4} is Abelian, A_{4}/V_{4} is cyclic of order 3, and Σ_{4}/A_{4} is cyclic of order 2.

Lemma 2.54 Let G be a group, let H_{1} and H_{2} be subgroups of G, where H_{1} ◁ H_{2}, and let J_{1} = H_{1} ∩ N, J_{2} = H_{2} ∩ N, K_{1} = H_{1}N/N and K_{2} = H_{2}N/N, where N is some normal subgroup of G. Then J_{1} ◁ J_{2} and K_{1} ◁ K_{2}. Moreover there exists a normal subgroup of H_{2}/H_{1} isomorphic to J_{2}/J_{1}, and the quotient of H_{2}/H_{1} by this normal subgroup is isomorphic to K_{2}/K_{1}.

Proof It is a straightforward exercise to verify that J_{1} ◁ J_{2} and K_{1} ◁ K_{2}. Let θ: H_{2} → K_{2} be the surjective homomorphism sending h ∈ H_{2} to the coset hN. Now θ induces a well-defined surjective homomorphism ψ: H_{2}/H_{1} → K_{2}/K_{1}, since θ(H_{1}) ⊂ K_{1}. Also θ^{−1}(K_{1}) = H_{2} ∩ (H_{1}N). But H_{2} ∩ (H_{1}N) = H_{1}(H_{2} ∩ N), for if a ∈ H_{1}, b ∈ N and ab ∈ H_{2} then b ∈ H_{2} ∩ N. Therefore

ker ψ = θ^{−1}(K_{1})/H_{1} = H_{1}(H_{2} ∩ N)/H_{1} ≅ H_{2} ∩ N/H_{1} ∩ N = J_{2}/J_{1}

by the First Isomorphism Theorem (Theorem 2.23). Moreover the quotient of H_{2}/H_{1} by the normal subgroup ker ψ is isomorphic to the image K_{2}/K_{1} of ψ. Thus ker ψ is the required normal subgroup of H_{2}/H_{1}.

Proposition 2.55 Let G be a group, and let H be a subgroup of G. Then

(i) if G is solvable then any subgroup H of G is solvable;

(ii) if G is solvable then G/N is solvable for any normal subgroup N of G;

(iii) if N is a normal subgroup of G and if both N and G/N are solvable then G is solvable.

Proof Suppose that G is solvable. Let G_{0}, G_{1}, …, G_{m} be a finite sequence of subgroups of G, where G_{0} = {1}, G_{n} = G, and G_{i−1} ◁ G_{i} and G_{i}/G_{i−1} is Abelian for i = 1, 2, …, m.

We first show that the subgroup H is solvable. Let H_{i} = H ∩ G_{i} for i = 0, 1, …, m. Then H_{0} = {1} and H_{m} = H. If u ∈ H_{i} and v ∈ H_{i−1} then uvu^{−1} ∈ H, since H is a subgroup of G. Also uvu^{−1} ∈ G_{i−1}, since u ∈ G_{i−1}, v ∈ G_{i} and G_{i−1} is normal in G_{i}. Therefore uvu^{−1} ∈ H_{i−1}. Thus H_{i−1} is a normal subgroup of H_{i} for i = 1, 2, …, m. Moreover

\frac{H_{i}}{H_{i−1}} = \frac{G_{i} ∩ H}{G_{i−1} ∩ (G_{i} ∩ H)} ≅ \frac{G_{i−1}(G_{i} ∩ H)}{G_{i−1}}

by the First Isomorphism Theorem (Theorem 2.23), and thus H_{i}/H_{i−1} is isomorphic to a subgroup of the Abelian group G_{i}/G_{i−1}. It follows that H_{i}/H_{i−1} must itself be an Abelian group. We conclude therefore that the subgroup H of G is solvable.

Now let N be a normal subgroup of G, and let K_{i} = G_{i}N/N for all i. Then K_{0} is the trivial subgroup of G/N and K_{m} = G/N. It follows from Lemma 2.54 that K_{i−1} ◁ K_{i} and K_{i}/K_{i−1} is isomorphic to the quotient of G_{i}/G_{i−1} by some normal subgroup. But a quotient of any Abelian group must itself be Abelian. Thus each quotient group K_{i}/K_{i−1} is Abelian, and thus G/N is solvable.

Finally suppose that G is a group, N is a normal subgroup of G and both N and G/N are solvable. We must prove that G is solvable. Now the solvability of N ensures the existence of a finite sequence G_{0}, G_{1}, …, G_{m} of subgroups of N, where G_{0} = {1}, G_{m} = N, and G_{i−1} ◁ G_{i} and G_{i}/G_{i−1} is Abelian for i = 1, 2, …, m. Also the solvability of G/N ensures the existence of a finite sequence K_{0}, K_{1}, …, K_{n} of subgroups of G/N, where K_{0} = N/N, K_{n} = G/N, and K_{i−1} ◁ K_{i} and K_{i}/K_{i−1} is Abelian for i = 1, 2, …, n. Let G_{m+i} be the preimage of K_{i} under the the quotient homomorphism ν: G → G/N, for i = 1, 2, …, n. The Second Isomorphism Theorem (Theorem 2.24) ensures that G_{m+i}/G_{m+i−1} ≅ K_{i}/K_{i−1} for all i > 0. Therefore G_{0}, G_{1}, …, G_{m+n} is a finite sequence of subgroups of G, where G_{0} = {1}, G_{n} = G, and G_{i−1} ◁ G_{i} and G_{i}/G_{i−1} is Abelian for i = 1, 2, …, m + n. Thus the group G is solvable, as required.

Example The alternating group A_{n} is simple for n ≥ 5 (see Theorem 2.36). Moreover the definition of solvable groups ensures that that any simple solvable group is cyclic, and A_{n} is not cyclic when n ≥ 5. Therefore A_{n} is not solvable when n ≥ 5. It then follows from Proposition 2.55 that the symmetric group Σ_{n} is not solvable when n ≥ 5.